### v1

# compat=1.1.0

Abstract
The Haldane model is a standard tight-binding model describing electrons hopping on a hexagonal lattice subject to a transverse, dipolar magnetic field. We consider its interacting version for values of the interaction strength that are small compared to the bandwidth. We study the critical case at the transition between the trivial and the ‘topological’ insulating phases, and we rigorously establish that the transverse conductivity on the dressed critical line is quantized at a half-integer multiple of $e^{2}/h$: this is the average of the integer values of the Hall conductivity in the insulating phases on either side of the dressed critical line. Together with previous results, this fully characterizes the nature of the phase transition between different Hall plateaus and proves its universality with respect to many-body interactions. The proof is based on a combination of constructive renormalization group methods and exact lattice Ward identities.
I Introduction
One of the central questions of solid state physics is the effect of disorder and interactions on quantum transport coefficients. A particularly interesting problem is to understand the dependence, if any, of the transverse conductivity of two-dimensional electron systems subject to an external, transverse magnetic field, on disorder and interactions. It is very well known [1, 2, 3, 4] that, in the independent electron approximation, the transverse conductivity of 2D lattice electron systems with or without disorder, and Fermi energy lying in a spectral or in a mobility gap, is quantized in integer multiples of $e^{2}/h$, a phenomenon known as the integer quantum Hall effect. In this setting, quantization follows from the observation that the Kubo conductivity is proportional to a geometrical index, the first Chern number associated with the ‘Bloch bundle’, or the Fredholm index of an appropriate pair of Fermi projectors [2, 3, 5].
For interacting systems, quantization in integer multiples of $e^{2}/h$ for gapped many-fermion systems follows from the interpretation of the Kubo conductivity in terms of a many-body geometric index [6, 7, 8]. This approach requires that the interacting system displays a finite spectral gap in the thermodynamic limit, which can be typically proved only for weak perturbations of gapped independent electron systems, where ‘weak’ means here that the interaction strength is much smaller than the non-interacting gap [9, 10].
On the other hand, a fundamental understanding of interaction effects on the transverse conductivity in systems that, in the absence of interactions, are gapless, is extremely challenging and, in most respects, still open. Two common and important settings where this question is relevant are: the fractional quantum Hall effect, which concerns electron systems subject to an external magnetic field at special fillings such that a gap is expected to open thanks to the interaction (mass generation at fractional fillings, a phenomenon that is mostly unexplained at a fundamental, microscopic level); and the critical phase corresponding to the transition from one integer quantum Hall plateau to another.
In this paper, we investigate the nature of the ‘topological transition’ from the normal insulating phase to a non-trivial quantum Hall phase in a specific class of 2D interacting electron systems, characterized by a critical semimetallic behavior at the transition. The approach we follow is not based on the introduction and use of geometrical indices or topological invariants. Rather, it combines the use of Ward identities and Schwinger–Dyson equations within a constructive, rigorous renormalization group (RG) scheme. The fact that quantization of the transport coefficients can be inferred from Ward identities and Schwinger–Dyson equations is not new [11, 12], and is related to the non-renormalization property of quantum anomalies [13, 14], as stated e.g. in the Adler–Bardeen theorem [15, 16]. Implementing these ideas within a constructive RG scheme is relatively new and, in our view, important, in that it allows one to unambiguously prove the universality, or non-universality, of transport coefficients, by fully taking into account finite effects due to irrelevant terms in the microscopic Hamiltonian: note that these are very difficult, if not impossible, to take into account within formal schemes based on an effective field theory description of the system, or on standard perturbative treatments. A remarkable case in which neglecting irrelevant terms associated with lattice effects leads to wrong predictions is that of the optical conductivity of graphene with short-range or Coulomb interactions, where different studies of effective models of interacting graphene based on ultraviolet-regularized interacting Dirac fermions led to contradictory results, in disagreement with the experiments [17, 18, 19, 20, 21]: in the case of short-range interactions, the use of constructive RG methods allowed to resolve these ambiguities and to rigorously prove the universality of the optical conductivity [22, 23]. These methods have also been used to rigorously prove the universality of transport coefficients of several other interacting Fermi systems in one, two and three dimensions, including the Drude weight of non-integrable quantum spin chains [24], the longitudinal conductivity of the Haldane–Hubbard model on the critical line [25, 26], and the condensed matter analogue of the chiral anomaly in Weyl semimetals [27].
A case that remained elusive so far is that of the transverse conductivity in semimetallic critical phases, such as those at the transition between different quantum Hall phases in 2D interacting electron systems on the hexagonal lattice (recall that systems with hexagonal symmetry generically display a semimetallic critical behavior [28]). The expectation, based on computations performed in non-interacting systems of Dirac fermions, is that in such a setting the critical transverse conductivity is quantized in a half-integer multiple of $e^{2}/h$, equal to the average of the two integer multiples displayed on the two different sides of critical state, see [29] and [30, Eq.(340)]. It is then argued that interactions cannot change this semi-integer value, because they are either explicitly irrelevant or marginally irrelevant as in the special case of Coulomb interactions; however, as discussed above in the context of the optical conductivity of graphene, this argument is inconclusive, because irrelevant terms can in general modify the values of finite quantities, including the conductivity, unless they are protected by symmetries. In this paper, we rigorously prove the quantization of the critical transverse conductivity in semi-integer multiples of $e^{2}/h$ in the setting of the Haldane–Hubbard model, i.e., the Haldane model [31] perturbed by a generalized Hubbard interaction, which we started to investigate in a series of previous papers [25, 32, 33, 26]. The restriction to this specific setting is done just for technical simplicity and not for any physically compelling reason. We expect that our proof extends to a wider class of 2D interacting Fermi systems with critical semimetallic behavior, but we postpone such a generalization to a future publication.
We recall that the Haldane model, in its non-interacting version, describes tight-binding electrons on the honeycomb lattice, subject to a transverse, dipolar magnetic field with zero net flux through the unit cell. The electrons can hop between nearest sites, with hopping strength $t_{1}$, and next-to-nearest sites, with hopping strength $t_{2}e^{i\phi}$ or $t_{2}e^{-i\phi}$, depending on the orientation of the next-to-nearest neighbor hopping (see Fig. 3), where $\phi$ represents the line integral between the two points of the vector potential generating the magnetic field. The electrons are also subject to a local staggered potential, which takes values $+W$ and $-W$ on the even and odd sublattices of the honeycomb lattice, respectively. Assuming that $30<t_{2}<t_{1}/$ 3, for generic values of $\phi,W$ the valence and conduction bands are separated by a spectral gap. However, there are two critical curves in the $(\phi,W)$ plane, $W=\pm 3\sqrt{3}t_{2}\sin\phi$, at which the two bands touch: they divide the plane $(\phi,W)$ in four disconnected regions (see Fig. 1), where the energy spectrum has a non-vanishing gap and the system exhibits an integer quantum Hall effect.
Figure 1: The conductivity matrix $\sigma$ in units of $\frac{e^{2}}{h}$ over the whole phase diagram (critical lines included). At the graphene points $(\phi,W)=(0,0),(\pi,0)$, the conductivity matrix is $\pi/4$ times the identity.
Two of these regions correspond to the topological insulator phase (TI), and are characterized by a non-trivial Hall conductance $\sigma_{12}=-\sigma_{21}=\pm\frac{e^{2}}{h}$, while the other two correspond to the normal insulator phase (NI), with vanishing Hall conductance. On the two critical lines, $W=\pm 3\sqrt{3}t_{2}\sin\phi$, the bands touch at a single point in the Brillouin zone, called the Fermi point (whose location depends on the choice of the sign in the equation of the critical line), around which the effective dispersion relation is approximately linear. The graphene points $(\phi,W)=(0,0),(\pi,0)$ are special, in that the bands touch simultaneously at both Fermi points, i.e., they display two conical intersections. The critical behavior of the system, associated with the transition between two plateaus of integer Hall conductivity, is, therefore, of semimetallic character. At the phase transition, both the longitudinal and transverse conductivities are quantized and non-trivial: in particular, the critical Hall conductivity turns out to be quantized at the half-integer multiple of $e^{2}/h$ equal to the average of the Hall conductivities on the two sides of the critical line. The computation of the critical conductivity matrix in the Haldane model is very amusing and instructive, but we could not find it in the literature: the closest computation we found is the one in [29, Appendix 1], which, however, is based on a relativistic, linear approximation of the Hamiltonian around the Fermi points. In the lattice setting, the result is implicitly stated in Haldane’s original paper, see [31, p.2017, column 2]: roughly speaking, the argument is that at the transition, one of the two Weyl components of the Dirac fermion is ‘heavy’ and breaks time reversal symmetry, while the other is massless; the Berry curvature is not distributed uniformly over the Brillouin zone but is highly concentrated near the band crossings: so when the mass of one Weyl fermion vanishes the other still has 1/2 of the Chern number, which implies that at the transition $\sigma_{12}=\pm\tfrac{1}{2}\tfrac{e^{2}}{h}$. A consistent computation for the full model, substantiating this argument, and performed without neglecting non-linear lattice effects, is presented in Section III.1 below.
The interacting version of the Haldane model we consider is obtained by adding a generalized, finite-range, Hubbard interaction of strength $U$. The chemical potential is fixed in the middle of the gap between the valence and conduction bands: at criticality, it is tuned precisely to the (renormalized) energy at the Fermi points. In a series of previous works, we proved the analyticity of the ground state static correlation functions, we computed the Kubo conductivity and proved its universality, i.e., its independence from the interaction strength, in various regimes. The easiest case to handle is when $(\phi,W)$ is fixed away from the non-interacting critical lines and the interaction strength $U$ is sufficiently small as compared to the unperturbed gap: in this case, ‘naive’ perturbation theory converges and a combined use of Ward identities and Schwinger–Dyson equations, in the spirit of [11, 12], allowed us to prove the quantization of the Hall conductivity [32, 33]. In order to extend the result to all non-critical values of the parameters, in particular to situations where $U$ is small compared to the bandwidth but in general larger than the unperturbed gap, one needs to apply a multiscale RG scheme, with which we derived the equation for the dressed critical lines and extended the universality of the transverse conductivity to all values of $(\phi,W)$ outside the dressed critical lines [26]. In this setting, we also succeeded in proving the universality of the longitudinal conductivity on the dressed critical lines [25].
As mentioned above, the case of the transverse conductivity on the dressed critical lines remained elusive. The reason why the critical transverse conductivity is more difficult to compute than the longitudinal one, is that while the latter is dominated by relativistic contributions (the irrelevant lattice contributions being zero by parity), the critical Hall conductivity is dominated by irrelevant contributions, i.e., by quasi-momenta away from the Fermi points. A priori, it is unclear how to evaluate them effectively in a multiscale computation: constructive RG typically allows one to isolate explicit, dominant, relativistic contributions from the subleading, irrelevant ones, which are finite and equal to the convergent sum of infinitely many Feynman diagrams, but there is no simple way to evaluate them explicitly. The key to the solution, described below, is the comparison of the critical value of the transverse conductivity with the average of its values at a distance $\epsilon$ inside and outside the curve: the difference between the critical value and such an average is now dominated by explicit relativistic contributions, which are shown to be zero; the irrelevant contributions to the difference are also zero, in the limit of sending the regularization parameter $\epsilon$ to zero (see the discussion after the statement of the main theorem in Section II.4 below for a more detailed description of this strategy). The resulting interacting topological phase diagram is shown in Fig.2, which summarizes the findings of this and previous papers, and fully determines the conductivity matrix of the model for all possible choices of $(\phi,W)$, provided that $U$ is sufficiently small compared to the bandwidth.
Figure 2: The phase diagram of the Haldane–Hubbard model with the corresponding conductivity matrix. The effect of the interaction shows in the renormalization of the critical curves (solid lines) compared to their non-interacting counterparts (dashed lines). According to [26], the critical curves are continuously differentiable in $\phi$ and $O(U)$-close to the non-interacting ones (see [25] for the leading-order computation). The conductivity matrix on and away from the dressed critical lines is the same as for the non-interacting model and, in particular, is $U$-independent.
This paper is organized as follows. In Sec.II, we define the model and the main observables to be analyzed in this work. In Sec.II.4 we state our main result, which is summarized in Theorem II.1. In Sec.III.1, we compute the transverse conductivity in the non-interacting Haldane model, while in Sec.III.2 we give the proof for the interacting case, under the assumption that the Kubo conductivity, as computed from the RG multiscale expansion, can be decomposed as a sum of two terms: a ‘relativistic’ contribution (see Eqs.(32)-(33)), which is dimensionally logarithmically divergent at criticality (but vanishing by parity reasons), plus a finite contribution from the irrelevant terms, which is more regular, i.e., continuous in the infrared cutoff (so that the difference between its value on the critical line and the average of its values on the two sides of the line vanishes as the infrared regularization parameter is sent to zero). The proof of the validity of this decomposition, which is summarized in Prop.III.1, relies on the RG strategy developed in [25, 26] and is discussed in Sec.IV.
II The model and the main results
II.1 Set-up
We shall think of the hexagonal lattice, which our model is defined on, as the superposition of two triangular lattices $\Lambda$ and $\Lambda^{\prime}$, shifted with respect to one another. We let $\Lambda=\cup_{n\in\mathbb{Z}^{2}}\{n_{1}\vec{\ell}_{1}+n_{2}\vec{\ell}_{2}\}$ be the infinite triangular lattice generated by the basis vectors
$\vec{\ell}_{1}=\frac{1}{2}(3,-\sqrt{3}),\hskip 28.45274pt\vec{\ell}_{2}=\frac{1}{2}(3,\sqrt{3}),$
and $\Lambda^{\prime}=\Lambda+(1,0)$. We then label each point of the hexagonal lattice by a pair $(\vec{x},\rho)$ with $\vec{x}\in\Lambda$ and $\rho\in\{1,2\}$, with the understanding that $(\vec{x},1)$ corresponds to $\vec{x}\in\Lambda$ and $(\vec{x},2)$ to $\vec{x}+(1,0)\in\Lambda^{\prime}$ (see Fig. 3). We denote by $\psi^{\dagger}_{\vec{x},\rho}$ and $\psi_{\vec{x},\rho}$ the fermionic creation and annihilation operators at site $(\vec{x},\rho)$, respectively, and by $\psi^{\dagger}_{\vec{x}}$ and $\psi_{\vec{x}}$ the corresponding $2$-component spinors.
Figure 3: The honeycomb lattice of the Haldane model, with the white (resp. black) points corresponding to the sublattice $\Lambda$ (resp. $\Lambda^{\prime}$). The nearest-neighbor vectors $\vec{\delta}_{1},\vec{\delta}_{2},\vec{\delta_{3}}$ and next-to-nearest neighbor vectors $\vec{\gamma}_{1},\vec{\gamma}_{2},\vec{\gamma}_{3}$ are also represented. For the latter, the phases associated to the hoppings from a white (resp. black) site to their next-to-nearest neighbor are shown in blue (resp. red); the hoppings in the opposite directions have complex conjugate phases.
Given a positive integer $L$, we study the model on the discrete torus $\Lambda_{L}=\Lambda/L\Lambda$; the box size $L$ acts as an infrared cutoff that will eventually be sent to infinity. We use the following convention for the Fourier transform of the fermionic operators:
$\hat{\psi}_{\vec{k}}=\sum_{\vec{x}\in\Lambda_{L}}e^{i\vec{k}\cdot\vec{x}}\psi_{\vec{x}},\hskip 14.22636pt\hat{\psi}_{\vec{x}}=\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}e^{-i\vec{k}\cdot\vec{x}}\psi_{\vec{k}},$
with $\vec{k}$ belonging to the discretized Brillouin zone
$\mathcal{B}_{L}:=\left\{\vec{k}=\tfrac{n_{1}}{L}\vec{G}_{1}+\tfrac{n_{2}}{L}\vec{G}_{2}:n_{i}\in\mathds{Z}\cap\left(-\tfrac{L}{2},\tfrac{L}{2}\right]\right\},$ (1)
which is defined in terms of the vectors $\vec{G}_{1},\vec{G}_{2}$ specified by the condition $\vec{G}_{i}\cdot\vec{\ell}_{j}=2\pi\delta_{ij}$. $\mathcal{B}_{L}$ should also be thought of as a discrete torus, i.e., the sum of any two vectors in $\mathcal{B}_{L}$ should be identified with an element of $\mathcal{B}_{L}$ modulo integer translations by $\vec{G}_{1}$ and/or $\vec{G}_{2}$. We let $\mathcal{B}\equiv\mathcal{B}_{\infty}$ be the infinite-volume Brillouin zone.
The non-interacting, finite-volume Hamiltonian with periodic boundary conditions is [31]
$\begin{split}&\mathcal{H}_{0,L}=-t_{1}\sum_{\vec{x}\in\Lambda_{L}}\left[\psi^{\dagger}_{\vec{x},1}\left(\psi_{\vec{x},2}+\psi_{\vec{x}-\vec{\ell}_{1},2}+\psi_{\vec{x}-\vec{\ell}_{2},2}\right)+h.c.\right]\\ &-t_{2}\sum_{\vec{x}\in\Lambda_{L}}\sum_{\begin{subarray}{c}\alpha=\pm\\ j=1,2,3\end{subarray}}\left(e^{i\alpha\phi}\psi^{\dagger}_{\vec{x},1}\psi_{\vec{x}+\alpha\vec{\gamma}_{j},1}+e^{-i\alpha\phi}\psi^{\dagger}_{\vec{x},2}\psi_{\vec{x}+\alpha\vec{\gamma}_{j},2}\right)\\ &+W\sum_{\vec{x}\in\Lambda_{L}}\left(\psi^{{\dagger}}_{\vec{x},1}\psi_{\vec{x},1}-\psi^{{\dagger}}_{\vec{x},2}\psi_{\vec{x},2}\right),\end{split}$
where: the first line describes nearest neighbor hopping by vectors $\vec{\delta}_{1}=(1,0)$, $\vec{\delta}_{2}=(-\tfrac{1}{2},\tfrac{\sqrt{3}}{2})$, $\vec{\delta}_{3}=(-\tfrac{1}{2},-\tfrac{\sqrt{3}}{2})$, see Fig.3; the second describes next-to-nearest neighbor hopping by vectors
$\vec{\gamma}_{1}=\vec{\ell}_{1}-\vec{\ell}_{2},\ \vec{\gamma}_{2}=\vec{\ell}_{2},\ \vec{\gamma}_{3}=-\vec{\ell}_{1},$
see again Fig.3; and the third describes a staggered local potential, favoring occupancy of the black (resp.  white) sites for $W>0$ (resp. $W<0$). We fix the hopping strengths $t_{1},t_{2}$ once and for all; for definiteness, we assume $t_{1},t_{2}>0$; moreover, in order to avoid having overlapping bands (see Section II.3), we assume that $t_{2}$ is not too large as compared to $t_{1}$, namely $t_{2}/t_{1}<1/3$. Given $t_{1},t_{2}$ with these properties, the non-interacting Hamiltonian will be thought of as being parametrized by $W$ and $\phi$, and in order to make such dependence explicit, we shall write $\mathcal{H}_{0,L}=\mathcal{H}_{0,L}(W,\phi)$.
The interacting Hamiltonian in the grand-canonical ensemble is
$\mathcal{H}_{L}=\mathcal{H}_{0,L}(W,\phi)-\mu\mathcal{N}_{L}+U\mathcal{V}_{L},$ (2)
which has to be thought of as a function of the parameters $W,\phi,\mu,U$, where $\mu$ is the chemical potential and $U$ is the interaction strength; in (2), letting $n_{\vec{x},\rho}=\psi^{\dagger}_{\vec{x},\rho}\psi_{\vec{x},\rho}$ be the local density, $\mathcal{N}_{L}=\sum_{\vec{x}\in\Lambda_{L}}\sum_{\rho=1,2}n_{\vec{x},\rho}$ is the total number operator, and $\mathcal{V}_{L}$ is a generalized Hubbard interaction, of the form
$\mathcal{V}_{L}=\sum_{\vec{x},\vec{y}\in\Lambda_{L}}\sum_{\rho,\rho^{\prime}\in\{1,2\}}(n_{\vec{x},\rho}-\tfrac{1}{2})v_{\rho,\rho^{\prime}}(\vec{x}-\vec{y})(n_{\vec{y},\rho^{\prime}}-\tfrac{1}{2}).$ (3)
The potential $v_{\rho,\rho^{\prime}}$ satisfies $v_{11}(\vec{x})=v_{22}(\vec{x})=v(\vec{x})$, $v_{12}(\vec{x})=v(\vec{x}-(1,0))$, $v_{21}(\vec{x})=v(\vec{x}+(1,0))$, with $v$ a rotationally-invariant, finite-range potential. The terms $-\tfrac{1}{2}$ in the definition of $\mathcal{V}_{L}$ provide a useful normalization of the chemical potential, which turns out to be convenient in the Grassmann representation of the model described below. In fact, thanks to the presence of the terms $-\tfrac{1}{2}$ in (3), the Grassmann counterpart of the interaction $\mathcal{V}_{L}$ is purely quartic, see the first line of (39) (for a proof that the Hamiltonian and Grassmann versions of the quartic interaction differ by these $-\tfrac{1}{2}$ factors, see [32, Sect.5.1], compare in particular [32, eq.(5.2)] and [32, eq.(5.9)]).
II.2 Correlation functions and Kubo conductivity
The central object of our interest is the Kubo conductivity, which is written in terms of the two-point current correlation function. We recall that the lattice current operator is defined by first promoting the Hamiltonian $\mathcal{H}_{L}$ to be $\vec{A}$-dependent, where $\vec{A}$ is an external $U(1)$ vector potential, via a replacement of the hopping parameters to $\vec{A}$-dependent hoppings via the Peierls substitution; and then by taking $\vec{\mathcal{J}}_{\vec{p}}=\delta H(\vec{A})/\delta\vec{A}_{\vec{p}}\big{|}_{\vec{A}=\vec{0}}$, see [25, Appendix A]; or, equivalently, it can be defined by imposing the lattice continuity equation, see [26, Sect.2.2]. The result is
$\vec{\mathcal{J}}_{\vec{p}}=\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}\hat{\psi}^{\dagger}_{\vec{k}+\vec{p}}\vec{\Gamma}(\vec{k},\vec{p})\hat{\psi}_{\vec{k}},$ (4)
where $\vec{p}\in\mathcal{D}_{L}:=\{\vec{k}=\tfrac{n_{1}}{L}\vec{G}_{1}+\tfrac{n_{2}}{L}\vec{G}_{2}:n_{i}\in\mathds{Z}\}$, and the components of the vector $\vec{\Gamma}(\vec{k},\vec{p})$ are $2\times 2$ matrices, called the bare vertex functions, such that
$\begin{split}&\vec{\Gamma}_{1,1}(\vec{k},\vec{p})=-it_{2}\sum_{j=1}^{3}\sum_{\alpha=\pm}\alpha\vec{\gamma}_{j}\eta_{\alpha\vec{\gamma}_{j}\cdot\vec{p}}e^{i\alpha(\phi-\vec{k}\cdot\vec{\gamma}_{j})}\\ &\vec{\Gamma}_{1,2}(\vec{k},\vec{p})=-it_{1}\sum_{j=1}^{3}\vec{\delta}_{j}\eta_{\vec{\delta}_{j}\cdot\vec{p}}e^{-i\vec{k}\cdot(\vec{\delta}_{j}-\vec{\delta}_{1})},\end{split}$ (5)
with $\eta_{x}:=(e^{-ix}-1)/(-ix)$, while $\vec{\Gamma}_{2,1}(\vec{k},\vec{p})=-\vec{\Gamma}_{1,2}(-\vec{k}-\vec{p},\vec{p})$, and $\vec{\Gamma}_{2,2}(\vec{k},\vec{p})=-e^{-i\vec{p}\cdot\vec{\delta}_{1}}\vec{\Gamma}_{1,1}(-\vec{k},-\vec{p})$.
In particular, the current operator at $\vec{p}=\vec{0}$ can be written as (cf. [32, Appendix A])
$\vec{\mathcal{J}}_{\vec{0}}=-\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}\tilde{\psi}^{\dagger}_{\vec{k}}\nabla_{\vec{k}}\tilde{H}^{0}(\vec{k})\tilde{\psi}_{\vec{k}},$ (6)
where $\tilde{\psi}_{\vec{k}}=\Gamma^{\dagger}_{0}(\vec{k})\hat{\psi}_{\vec{k}}$ and $\tilde{H}^{0}(\vec{k})=\Gamma^{\dagger}_{0}(\vec{k})\hat{H}^{0}(\vec{k})\Gamma_{0}(\vec{k})$. Here, $\Gamma_{0}(\vec{k})=\begin{pmatrix}1&0\\ 0&e^{-ik_{1}}\end{pmatrix}$ and $\hat{H}^{0}(\vec{k})$ is the Bloch Hamiltonian of the non-interacting system (cf. [31] and [32, Appendix B]):
$\hat{H}^{0}(\vec{k})=$ (7)
$\left(\begin{array}[]{cc}-2t_{2}\alpha_{1}(\vec{k})\cos\phi+m(\vec{k})&-t_{1}\Omega^{*}(\vec{k})\\ -t_{1}\Omega(\vec{k})&-2t_{2}\alpha_{1}(\vec{k})\cos\phi-m(\vec{k})\\ \end{array}\right),$ (10)
where
$\alpha_{1}(\vec{k})$ $=\sum_{j=1}^{3}\cos(\vec{k}\cdot\vec{\gamma}_{j}),\hskip 14.22636pt$ $m(\vec{k})=W-2t_{2}\alpha_{2}(\vec{k})\sin\phi,$
$\alpha_{2}(\vec{k})$ $=\sum_{j=1}^{3}\sin(\vec{k}\cdot\vec{\gamma}_{j}),\hskip 14.22636pt$ $\Omega(\vec{k})=1+e^{-i\vec{k}\cdot\vec{\ell}_{1}}+e^{-i\vec{k}\cdot\vec{\ell}_{2}}.$
In the thermodynamic and zero-temperature limits, after a Wick rotation of the time variable, the Kubo conductivity matrix of elements $\sigma_{ij}$ can be expressed in terms of Euclidean current-current correlations, see [32, Theorem 3.1] for a proof of the validity of the Wick rotation for the conductivity in the off-critical case, and [34, Sec.5] for a proof in the critical case. One finds that
$\sigma_{ij}=\frac{1}{|\vec{\ell}_{1}\times\vec{\ell}_{2}|}\partial^{-}\hat{K}_{ij}(0),$ (11)
where $i,j\in\{1,2\}$, $|\vec{\ell}_{1}\times\vec{\ell}_{2}|=3\sqrt{3}/2$ is the unit cell area, $\partial^{-}\hat{K}_{ij}(0)=\lim_{p_{0}\to 0^{-}}\tfrac{\hat{K}_{ij}(p_{0})-\hat{K}_{ij}(0)}{p_{0}}$ is the left derivative at $p_{0}=0$, and
$\begin{split}&\hat{K}_{ij}(p_{0})=\lim_{\beta\to\infty}\lim_{L\to\infty}\frac{1}{\beta L^{2}}\int_{0}^{\beta}dt_{1}\int_{0}^{\beta}dt_{2}\,e^{-ip_{0,\beta}(t_{1}-t_{2})}\cdot\\ &\quad\cdot\Big{[}\langle T\big{(}\mathcal{J}_{\vec{0},i}(t_{1})\mathcal{J}_{\vec{0},j}(t_{2})\big{)}\rangle_{\beta,L}-\langle\mathcal{J}_{\vec{0},i}\rangle_{\beta,L}\langle\mathcal{J}_{-\vec{0},j}\rangle_{\beta,L}\Big{]}\end{split}$ (12)
is the Euclidean current-current correlation at Matsubara frequency $p_{0}$ in the thermodynamic and zero-temperature limits. In (12), the expectation $\langle\cdot\rangle_{\beta,L}$ is computed with respect to the Gibbs measure at inverse temperature $\beta$,
$\langle\cdot\rangle_{\beta,L}=\frac{\mbox{Tr}\left\{e^{-\beta\mathcal{H}_{L}}\cdot\right\}}{\mbox{Tr}\left\{e^{-\beta\mathcal{H}_{L}}\right\}},$
and, for any $0\leq t<\beta$, $\mathcal{J}_{\vec{p},i}(t)=e^{t\mathcal{H}_{L}}\mathcal{J}_{\vec{p},i}e^{-t\mathcal{H}_{L}}$ is the imaginary time evolution of the current; moreover, $p_{0,\beta}=\tfrac{2\pi}{\beta}\lfloor\tfrac{\beta p_{0}}{2\pi}\rfloor$ is an integer multiple of $2\pi/\beta$ tending to $p_{0}$ as $\beta\to\infty$; finally, the operator $T$ inside the expectation on the RHS is the time-ordering operator, which reorders the product of the two time-dependent operators in its argument in decreasing time order.
For later reference, we also introduce the two-point Schwinger function in the thermodynamic and zero-temperature limits (i.e., the Euclidean Green’s function)
$\begin{split}\hat{S}(k_{0},\vec{k})&=\lim_{\beta\to\infty}\lim_{L\to\infty}\tfrac{1}{\beta}\int_{0}^{\beta}dt_{1}\int_{0}^{\beta}dt_{2}\sum_{\vec{x}\in\Lambda_{L}}\cdot\\ &\cdot e^{ik_{0,\beta}(t_{1}-t_{2})+i\vec{k}_{L}\cdot\vec{x}}\langle T\big{(}\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big{)}\rangle_{\beta,L},\end{split}$ (13)
where $k_{0,\beta}=\tfrac{2\pi}{\beta}(\lfloor\tfrac{\beta\omega}{2\pi}\rfloor+\tfrac{1}{2})$, $(\vec{k}_{L})_{i}=\tfrac{2\pi}{L}\lfloor\tfrac{Lk_{i}}{2\pi}\rfloor$, and, again, $\psi^{(\dagger)}_{\vec{x}}(t)=e^{t\mathcal{H}_{L}}\psi^{(\dagger)}_{\vec{x}}e^{-t\mathcal{H}_{L}}$ indicates the time evolution of the creation or annihilation fermionic operators, while $T$ is the fermionic time ordering operator such that
$T\big{(}\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big{)}=\begin{cases}\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})&\text{if $t_{1}>t_{2}$}\\ -\psi^{\dagger}_{\vec{0}}(t_{2})\psi_{\vec{x}}(t_{1})&\text{if $t_{1}\leq t_{2}$}.\end{cases}$
II.3 The non-interacting theory
Before proceeding further, it is convenient to briefly recall a few properties of the system at $U=0$, in which case the model is exactly solvable. The energy bands, i.e., the $\vec{k}$-dependent eigenvalues of the Bloch Hamiltonian (7), are
$\varepsilon_{\pm}(\vec{k})=-2t_{2}\alpha_{1}(\vec{k})\cos\phi\pm\sqrt{m(\vec{k})^{2}+t_{1}^{2}|\Omega(\vec{k})|^{2}}.$
Under the assumption $t_{2}/t_{1}<1/3$, the two bands do not overlap, and can only touch at the two Fermi points, where $\Omega(\vec{k})$ vanishes: $\vec{k}_{F}^{\,\pm}=\frac{2\pi}{3}\big{(}1,\pm\tfrac{1}{\sqrt{3}}\big{)}$. Denoting $m_{\omega}:=m(\vec{k}_{F}^{\omega})$, with $\omega\in\{+,-\}$, the two bands touch at $\vec{k}_{F}^{\omega}$ iff
$m_{\omega}(\phi,W)=W+3\sqrt{3}\omega t_{2}\sin\phi=0.$ (14)
This equation defines two curves in the $(\phi,W)$-plane, which we call the critical curves of the non-interacting theory. We fix the chemical potential at $\mu=-2t_{2}\alpha_{1}(\vec{k}_{F}^{+})\cos\phi=3t_{2}\cos\phi$, so that, if $(\phi,W)$ is on the critical curves, the system is a semimetal, while, in the complement of the critical curves, it is in an insulating (gapped) phase. In that case, the computation of the conductivity matrix leads to the values indicated in Fig. 1: the computation in the off-critical, insulating case was discussed in [32, Appendix B]; the computation of the longitudinal conductivity on the critical lines was discussed in [22, Sect.IV] (at the graphene point) and in [25, Sect.IV.B] (away from the graphene point); the computation of the transverse conductivity on the criticial lines, away from the graphene points where it is trivially zero, is discussed in Section III.1 below.
For completeness, let us conclude this section by describing the form that the two-point Schwinger function (13) takes in the non-interacting case,
$\hat{S}^{0}(k_{0},\vec{k}):=\hat{S}(k_{0},\vec{k})\big{|}_{U=0}=\big{(}-ik_{0}+\hat{H}^{0}(\vec{k})-\mu\big{)}^{-1},$ (15)
with $\mu=3t_{2}\cos\phi$. Its Fourier dual is denoted by $S^{0}(t_{1}-t_{2},\vec{x})=\lim_{\beta\to\infty}\lim_{L\to\infty}\langle T\big{(}\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big{)}\rangle_{\beta,L}\big{|}_{U=0}$, and its partial Fourier dual (with respect to the Matsubara frequency) by $\widetilde{S}^{0}(t,\vec{k}):=\sum_{\vec{x}\in\Lambda}e^{i\vec{k}\cdot\vec{x}}S^{0}(t,\vec{x})$, which reads
$\tilde{S}^{0}(t,\vec{k})=e^{-t(\hat{H}^{0}(\vec{k})-\mu)}\left(\mathds{1}_{\{t>0\}}P_{+}(\vec{k})-\mathds{1}_{\{t\leq 0\}}P_{-}(\vec{k})\right),$ (16)
where $P_{+}(\vec{k})$ and $P_{-}(\vec{k})$ are the projections over the upper and lower bands, respectively.
If both $m_{+}$ and $m_{-}$ are different from zero, then $\hat{S}^{0}(\mathbf{k})$ is real analytic in $\mathbf{k}\equiv(k_{0},\vec{k})$ over $\mathbb{R}\times\mathcal{B}$, and its Fourier dual $S^{0}(\mathbf{x})$ decays exponentially over ×$\mathbb{R}\times\Lambda$. Conversely, on the critical lines defined by (14), $\hat{S}^{0}(\mathbf{k})$ in (15) has a simple pole at $\mathbf{k}_{F}^{\omega}\equiv(0,\vec{k}_{F}^{\omega})$:
$\begin{split}&\hat{S}^{0}\left(\mathbf{k}_{F}^{\omega}+\mathbf{k}^{\prime}\right)=\\ &\left(\begin{array}[]{cc}-ik_{0}^{\prime}&\frac{3}{2}t_{1}(ik_{1}^{\prime}-\omega k_{2}^{\prime})\\ -\frac{3}{2}t_{1}(ik_{1}^{\prime}+\omega k_{2}^{\prime})&-ik_{0}^{\prime}\\ \end{array}\right)^{-1}\left(1+O(|\mathbf{k}^{\prime}|)\right),\end{split}$ (17)