### v3

# Universality of the topological phase transition in the interacting Haldane model

Abstract
The Haldane model is a standard tight-binding model describing electrons hopping on a hexagonal lattice subject to a transverse, dipolar magnetic field. We consider its interacting version for values of the interaction strength that are small compared to the bandwidth. We study the critical case at the transition between the trivial and the ‘topological’ insulating phases, and we rigorously establish that the transverse conductivity on the dressed critical line is quantized at a half-integer multiple of $e^{2}/h$: this is the average of the integer values of the Hall conductivity in the insulating phases on either side of the dressed critical line. Together with previous results, this fully characterizes the nature of the phase transition between different Hall plateaus and proves its universality with respect to many-body interactions. The proof is based on a combination of constructive renormalization group methods and exact lattice Ward identities.
I Introduction
One of the central questions of solid state physics is the effect of disorder and interactions on quantum transport coefficients. A particularly interesting problem is to understand the dependence, if any, of the transverse conductivity of two-dimensional electron systems subject to an external, transverse magnetic field, on disorder and interactions. It is very well known [1, 2, 3, 4] that, in the independent electron approximation, the transverse conductivity of 2D lattice electron systems with or without disorder, and Fermi energy lying in a spectral or in a mobility gap, is quantized in integer multiples of $e^{2}/h$, a phenomenon known as the integer quantum Hall effect. In this setting, quantization follows from the observation that the Kubo conductivity is proportional to a geometrical index, the first Chern number associated with the ‘Bloch bundle’, or the Fredholm index of an appropriate pair of Fermi projectors [2, 3, 5].
For interacting systems, quantization in integer multiples of $e^{2}/h$ for gapped many-fermion systems follows from the interpretation of the Kubo conductivity in terms of a many-body geometric index [6, 7, 8]. See also [9, 10] and references therein for the definition of a topological invariant in terms of interacting Green functions. These approaches require that the interacting system displays a finite spectral gap in the thermodynamic limit, which can be typically proved only for weak perturbations of gapped independent electron systems, where ‘weak’ means here that the interaction strength is much smaller than the non-interacting gap [11, 12].
On the other hand, a fundamental understanding of interaction effects on the transverse conductivity in systems that, in the absence of interactions, are gapless, is extremely challenging and, in most respects, still open. Two common and important settings where this question is relevant are: the fractional quantum Hall effect, which concerns electron systems subject to an external magnetic field at special fillings such that a gap is expected to open thanks to the interaction (mass generation at fractional fillings, a phenomenon that is mostly unexplained at a fundamental, microscopic level); and the critical phase corresponding to the transition from one integer quantum Hall plateau to another.
In this paper, we investigate the nature of the ‘topological transition’ from the normal insulating phase to a non-trivial quantum Hall phase in a specific class of 2D interacting electron systems, characterized by a critical semimetallic behavior, which is the generic one on the transition line separating two distinct topological phases, both in two and three dimensions [13, 14]. The approach we follow is not based on the introduction and use of geometrical indices or topological invariants. It is unclear whether this is at all possible for interacting semimetallic systems (see [15] for a topological interpretation of non-interacting semimetals). Therefore, rather than characterizing the Hall conductivity in terms of a geometrical index, we use a strategy that combines the use of Ward identities and Schwinger–Dyson equations within a constructive, rigorous renormalization group (RG) scheme (for an alternative RG approach to topological phase transitions, see also [16]). The fact that quantization of the transport coefficients can be inferred from Ward identities and Schwinger–Dyson equations is not new [17, 18], and is related to the non-renormalization property of quantum anomalies [19, 20, 21], as stated e.g. in the Adler–Bardeen theorem [22, 23].
Implementing these ideas within a constructive RG scheme is relatively new and, in our view, important, in that it allows one to unambiguously prove the universality, or non-universality, of transport coefficients, by fully taking into account finite effects due to irrelevant terms in the microscopic Hamiltonian: note that these are very difficult, if not impossible, to take into account within formal schemes based on an effective field theory description of the system, or on standard perturbative treatments. A remarkable case in which neglecting irrelevant terms associated with lattice effects leads to wrong predictions is that of the optical conductivity of graphene with short-range or Coulomb interactions, where different studies of effective models of interacting graphene based on ultraviolet-regularized interacting Dirac fermions led to contradictory results, in disagreement with the experiments [24, 25, 26, 27, 28]: in the case of short-range interactions, the use of constructive RG methods allowed to resolve these ambiguities and to rigorously prove the universality of the optical conductivity [29, 30]. These methods have also been used to rigorously prove the universality of transport coefficients of several other interacting Fermi systems in one, two and three dimensions, including the Drude weight of non-integrable quantum spin chains [31], the longitudinal conductivity of the Haldane–Hubbard model on the critical line [32, 33], and the condensed matter analogue of the chiral anomaly in Weyl semimetals [34].
A case that remained elusive so far is that of the transverse conductivity in semimetallic critical phases, such as those at the transition between different quantum Hall phases in 2D interacting electron systems on the hexagonal lattice (recall that systems with hexagonal symmetry generically display a semimetallic critical behavior [35]). The expectation, based on computations performed in non-interacting systems of Dirac fermions, is that in such a setting the critical transverse conductivity is quantized in a half-integer multiple of $e^{2}/h$, equal to the average of the two integer multiples displayed on the two different sides of critical state, see [36] and [37, Eq.(340)]. It is then argued that interactions cannot change this semi-integer value, because they are either explicitly irrelevant or marginally irrelevant as in the special case of Coulomb interactions; however, as discussed above in the context of the optical conductivity of graphene, this argument is inconclusive, because irrelevant terms can in general modify the values of finite quantities, including the conductivity, unless they are protected by symmetries. In this paper, we rigorously prove the quantization of the critical transverse conductivity in semi-integer multiples of $e^{2}/h$ in the setting of the Haldane–Hubbard model, i.e., the Haldane model [38] perturbed by a generalized Hubbard interaction, which we started to investigate in a series of previous papers [32, 39, 40, 33]. The restriction to this specific setting is done just for technical simplicity and not for any physically compelling reason. We expect that our proof extends to a wider class of 2D interacting Fermi systems with critical semimetallic behavior, but we postpone such a generalization to a future publication.
We recall that the Haldane model, in its non-interacting version, describes tight-binding electrons on the honeycomb lattice, subject to a transverse, dipolar magnetic field with zero net flux through the unit cell. The electrons can hop between nearest sites, with hopping strength $t_{1}$, and next-to-nearest sites, with hopping strength $t_{2}e^{i\phi}$ or $t_{2}e^{-i\phi}$, depending on the orientation of the next-to-nearest neighbor hopping (see Fig. 3), where $\phi$ represents the line integral between the two points of the vector potential generating the magnetic field. The electrons are also subject to a local staggered potential, which takes values +W+W and W-W on the even and odd sublattices of the honeycomb lattice, respectively. Assuming that $0<t_{2}<t_{1}/3$, for generic values of $\phi,W$ the valence and conduction bands are separated by a spectral gap. However, there are two critical curves in the $(\phi,W)$ plane, $W=\pm 3\sqrt{3}t_{2}\sin\phi$, at which the two bands touch: they divide the plane $(\phi,W)$ in four disconnected regions (see Fig. 1), where the energy spectrum has a non-vanishing gap and the system exhibits an integer quantum Hall effect.
Figure 1: The conductivity matrix $\sigma$ in units of $\frac{e^{2}}{h}$ over the whole phase diagram (critical lines included). At the graphene points $(\phi,W)=(0,0),(\pi,0)$, the conductivity matrix is $\pi/4$ times the identity.
Two of these regions correspond to the topological insulator phase (TI), and are characterized by a non-trivial Hall conductance $\sigma_{12}=-\sigma_{21}=\pm\frac{e^{2}}{h}$, while the other two correspond to the normal insulator phase (NI), with vanishing Hall conductance. On the two critical lines, $W=\pm 3\sqrt{3}t_{2}\sin\phi$, the bands touch at a single point in the Brillouin zone, called the Fermi point (whose location depends on the choice of the sign in the equation of the critical line), around which the effective dispersion relation is approximately linear. The graphene points $(\phi,W)=(0,0),(\pi,0)$ are special, in that the bands touch simultaneously at both Fermi points, i.e., they display two conical intersections. The critical behavior of the system, associated with the transition between two plateaus of integer Hall conductivity, is, therefore, of semimetallic character. At the phase transition, both the longitudinal and transverse conductivities are quantized and non-trivial: in particular, the critical Hall conductivity turns out to be quantized at the half-integer multiple of $e^{2}/h$ equal to the average of the Hall conductivities on the two sides of the critical line. The computation of the critical conductivity matrix in the Haldane model is very amusing and instructive, but we could not find it in the literature: the closest computation we found is the one in [36, Appendix 1], which, however, is based on a relativistic, linear approximation of the Hamiltonian around the Fermi points. In the lattice setting, the result is implicitly stated in Haldane’s original paper, see [38, p.2017, column 2]: roughly speaking, the argument is that at the transition one of the two Weyl components of the Dirac fermion is ‘heavy’ and breaks time reversal symmetry, while the other is massless; the Berry curvature is not distributed uniformly over the Brillouin zone but is highly concentrated near the band crossings: so when the mass of one Weyl fermion vanishes the other still has 1/2 of the Chern number, which implies that at the transition $\sigma_{12}=\pm\tfrac{1}{2}\tfrac{e^{2}}{h}$. A consistent computation for the full model, substantiating this argument, and performed without neglecting non-linear lattice effects, is presented in Section III.1 below.
The interacting version of the Haldane model we consider is obtained by adding a generalized, finite-range, Hubbard interaction of strength $U$. The chemical potential is fixed in the middle of the gap between the valence and conduction bands: at criticality, it is tuned precisely to the (renormalized) energy at the Fermi points. In a series of previous works, we proved the analyticity of the ground state static correlation functions, we computed the Kubo conductivity and proved its universality, i.e., its independence from the interaction strength, in various regimes. The easiest case to handle is when $(\phi,W)$ is fixed away from the non-interacting critical lines and the interaction strength $U$ is sufficiently small as compared to the unperturbed gap: in this case, ‘naive’ perturbation theory converges and a combined use of Ward identities and Schwinger–Dyson equations, in the spirit of [17, 18], allowed us to prove the quantization of the Hall conductivity [39, 40]. In order to extend the result to all non-critical values of the parameters, in particular to situations where $U$ is small compared to the bandwidth but in general larger than the unperturbed gap, one needs to apply a multiscale RG scheme, with which we derived the equation for the dressed critical lines and extended the universality of the transverse conductivity to all values of $(\phi,W)$ outside the dressed critical lines [33]. In this setting, we also succeeded in proving the universality of the longitudinal conductivity on the dressed critical lines [32].
As mentioned above, the case of the transverse conductivity on the dressed critical lines remained elusive. The reason why the critical transverse conductivity is more difficult to compute than the longitudinal one, is that while the latter is dominated by relativistic contributions (the irrelevant lattice contributions being zero by parity), the critical Hall conductivity is dominated by irrelevant contributions, i.e., by quasi-momenta away from the Fermi points. A priori, it is unclear how to evaluate them effectively in a multiscale computation: constructive RG typically allows one to isolate explicit, dominant, relativistic contributions from the subleading, irrelevant ones, which are finite and equal to the convergent sum of infinitely many Feynman diagrams, but there is no simple way to evaluate them explicitly. The key to the solution, described below, is the comparison of the critical value of the transverse conductivity with the arithmetic average of its values at a distance $\epsilon$ inside and outside the curve: the difference between the critical value and such an average is now dominated by explicit relativistic contributions, which are shown to be zero; the irrelevant contributions to the difference are also zero, in the limit of sending the regularization parameter $\epsilon$ to zero (see the discussion after the statement of the main theorem in Section II.4 below for a more detailed description of this strategy). The fact that the arithmetic, rather than some other weighted, average is the correct one to use comes from an emergent parity symmetry close to the transition line, as discussed in Remark 1 below. The resulting interacting topological phase diagram is shown in Fig.2, which summarizes the findings of this and previous papers, and fully determines the conductivity matrix of the model for all possible choices of $(\phi,W)$, provided that $U$ is sufficiently small compared to the bandwidth.
Figure 2: The phase diagram of the Haldane–Hubbard model with the corresponding conductivity matrix. The effect of the interaction shows in the renormalization of the critical curves (solid lines) compared to their non-interacting counterparts (dashed lines). According to [33], the critical curves are continuously differentiable in $\phi$ and $O(U)$-close to the non-interacting ones (see [32] for the leading-order computation). The conductivity matrix on and away from the dressed critical lines is the same as for the non-interacting model and, in particular, is $U$-independent.
This paper is organized as follows. In Sec.II, we define the model and the main observables to be analyzed in this work. In Sec.II.4 we state our main result, which is summarized in Theorem II.1. In Sec.III.1, we compute the transverse conductivity in the non-interacting Haldane model, while in Sec.III.2 we give the proof for the interacting case, under the assumption that the Kubo conductivity, as computed from the RG multiscale expansion, can be decomposed as a sum of two terms: a ‘relativistic’ contribution (see Eqs.(32)-(33)), which is dimensionally logarithmically divergent at criticality (but vanishing by parity reasons), plus a finite contribution from the irrelevant terms, which is more regular, i.e., continuous in the infrared cutoff (so that the difference between its value on the critical line and the average of its values on the two sides of the line vanishes as the infrared regularization parameter is sent to zero). The proof of the validity of this decomposition, which is summarized in Prop.III.1, relies on the RG strategy developed in [32, 33] and is discussed in Sec.IV.
II The model and the main results
II.1 Set-up
We shall think of the hexagonal lattice, which our model is defined on, as the superposition of two triangular lattices $\Lambda$ and $\Lambda^{\prime}$, shifted with respect to one another. We let $\Lambda=\cup_{n\in\mathbb{Z}^{2}}\{n_{1}\vec{\ell}_{1}+n_{2}\vec{\ell}_{2}\}$ be the infinite triangular lattice generated by the basis vectors
$\vec{\ell}_{1}=\frac{1}{2}(3,-\sqrt{3}),\hskip 28.45274pt\vec{\ell}_{2}=\frac{1}{2}(3,\sqrt{3}),$
and $\Lambda^{\prime}=\Lambda+(1,0)$. We then label each point of the hexagonal lattice by a pair $(\vec{x},\rho)$ with $\vec{x}\in\Lambda$ and $\rho\in\{1,2\}$, with the understanding that $(\vec{x},1)$ corresponds to $\vec{x}\in\Lambda$ and $(\vec{x},2)$ to $\vec{x}+(1,0)\in\Lambda^{\prime}$ (see Fig. 3). We denote by $\psi^{\dagger}_{\vec{x},\rho}$ and $\psi_{\vec{x},\rho}$ the fermionic creation and annihilation operators at site $(\vec{x},\rho)$, respectively, and by $\psi^{\dagger}_{\vec{x}}$ and $\psi_{\vec{x}}$ the corresponding $2$-component spinors.
Figure 3: The honeycomb lattice of the Haldane model, with the white (resp. black) points corresponding to the sublattice $\Lambda$ (resp. $\Lambda^{\prime}$). The nearest-neighbor vectors $\vec{\delta}_{1},\vec{\delta}_{2},\vec{\delta_{3}}$ and next-to-nearest neighbor vectors $\vec{\gamma}_{1},\vec{\gamma}_{2},\vec{\gamma}_{3}$ are also represented. For the latter, the phases associated to the hoppings from a white (resp. black) site to their next-to-nearest neighbor are shown in blue (resp. red); the hoppings in the opposite directions have complex conjugate phases.
Given a positive integer $L$, we study the model on the discrete torus $\Lambda_{L}=\Lambda/L\Lambda$; the box size $L$ acts as an infrared cutoff that will eventually be sent to infinity. We use the following convention for the Fourier transform of the fermionic operators:
$\hat{\psi}_{\vec{k}}=\sum_{\vec{x}\in\Lambda_{L}}e^{i\vec{k}\cdot\vec{x}}\psi_{\vec{x}},\hskip 14.22636pt{\psi}_{\vec{x}}=\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}e^{-i\vec{k}\cdot\vec{x}}\hat{\psi}_{\vec{k}},$
with $\vec{k}$ belonging to the discretized Brillouin zone
$\mathcal{B}_{L}:=\left\{\vec{k}=\tfrac{n_{1}}{L}\vec{G}_{1}+\tfrac{n_{2}}{L}\vec{G}_{2}:n_{i}\in\mathds{Z}\cap\left(-\tfrac{L}{2},\tfrac{L}{2}\right]\right\},$ (1)
which is defined in terms of the vectors $\vec{G}_{1},\vec{G}_{2}$ specified by the condition $\vec{G}_{i}\cdot\vec{\ell}_{j}=2\pi\delta_{ij}$. $\mathcal{B}_{L}$ should also be thought of as a discrete torus, i.e., the sum of any two vectors in $\mathcal{B}_{L}$ should be identified with an element of $\mathcal{B}_{L}$ modulo integer translations by $\vec{G}_{1}$ and/or $\vec{G}_{2}$. We let $\mathcal{B}\equiv\mathcal{B}_{\infty}$ be the infinite-volume Brillouin zone.
The non-interacting, finite-volume Hamiltonian with periodic boundary conditions is [38]
$\begin{split}&\mathcal{H}_{0,L}=-t_{1}\sum_{\vec{x}\in\Lambda_{L}}\left[\psi^{\dagger}_{\vec{x},1}\left(\psi_{\vec{x},2}+\psi_{\vec{x}-\vec{\ell}_{1},2}+\psi_{\vec{x}-\vec{\ell}_{2},2}\right)+h.c.\right]\\ &-t_{2}\sum_{\vec{x}\in\Lambda_{L}}\sum_{\begin{subarray}{c}\alpha=\pm\\ j=1,2,3\end{subarray}}\left(e^{i\alpha\phi}\psi^{\dagger}_{\vec{x},1}\psi_{\vec{x}+\alpha\vec{\gamma}_{j},1}+e^{-i\alpha\phi}\psi^{\dagger}_{\vec{x},2}\psi_{\vec{x}+\alpha\vec{\gamma}_{j},2}\right)\\ &+W\sum_{\vec{x}\in\Lambda_{L}}\left(\psi^{{\dagger}}_{\vec{x},1}\psi_{\vec{x},1}-\psi^{{\dagger}}_{\vec{x},2}\psi_{\vec{x},2}\right),\end{split}$
where: the first line describes nearest neighbor hopping by vectors $\vec{\delta}_{1}=(1,0)$, $\vec{\delta}_{2}=(-\tfrac{1}{2},\tfrac{\sqrt{3}}{2})$, $\vec{\delta}_{3}=(-\tfrac{1}{2},-\tfrac{\sqrt{3}}{2})$, see Fig.3; the second describes next-to-nearest neighbor hopping by vectors
$\vec{\gamma}_{1}=\vec{\ell}_{1}-\vec{\ell}_{2},\ \vec{\gamma}_{2}=\vec{\ell}_{2},\ \vec{\gamma}_{3}=-\vec{\ell}_{1},$
see again Fig.3; and the third describes a staggered local potential, favoring occupancy of the black (resp. white) sites for $W>0$ (resp. $W<0$). We fix the hopping strengths $t_{1},t_{2}$ once and for all; for definiteness, we assume $t_{1},t_{2}>0$; moreover, in order to avoid having overlapping bands (see Section II.3), we assume that $t_{2}$ is not too large as compared to $t_{1}$, namely $t_{2}/t_{1}<1/3$. Given $t_{1},t_{2}$ with these properties, the non-interacting Hamiltonian will be thought of as being parametrized by $W$ and $\phi$, and in order to make such dependence explicit, we shall write $\mathcal{H}_{0,L}=\mathcal{H}_{0,L}(W,\phi)$.
The interacting Hamiltonian in the grand-canonical ensemble is
$\mathcal{H}_{L}=\mathcal{H}_{0,L}(W,\phi)-\mu\mathcal{N}_{L}+U\mathcal{V}_{L},$ (2)
which has to be thought of as a function of the parameters $W,\phi,\mu,U$, where $\mu$ is the chemical potential and $U$ is the interaction strength; in (2), letting $n_{\vec{x},\rho}=\psi^{\dagger}_{\vec{x},\rho}\psi_{\vec{x},\rho}$ be the local density, $\mathcal{N}_{L}=\sum_{\vec{x}\in\Lambda_{L}}\sum_{\rho=1,2}n_{\vec{x},\rho}$ is the total number operator, and $\mathcal{V}_{L}$ is a generalized Hubbard interaction, of the form
$\mathcal{V}_{L}=\sum_{\vec{x},\vec{y}\in\Lambda_{L}}\sum_{\rho,\rho^{\prime}\in\{1,2\}}(n_{\vec{x},\rho}-\tfrac{1}{2})v_{\rho,\rho^{\prime}}(\vec{x}-\vec{y})(n_{\vec{y},\rho^{\prime}}-\tfrac{1}{2}).$ (3)
The potential $v_{\rho,\rho^{\prime}}$ satisfies $v_{11}(\vec{x})=v_{22}(\vec{x})=v(\vec{x})$, $v_{12}(\vec{x})=v(\vec{x}-(1,0))$, $v_{21}(\vec{x})=v(\vec{x}+(1,0))$, with $v$ a rotationally-invariant, finite-range potential. The terms $\tfrac{1}{2}$ in the definition of $\mathcal{V}_{L}$ provide a useful normalization of the chemical potential, which turns out to be convenient in the Grassmann representation of the model described below. In fact, thanks to the presence of the terms $\tfrac{1}{2}$ in (3), the Grassmann counterpart of the interaction $\mathcal{V}_{L}$ is purely quartic, see the first line of (39) (for a proof that the Hamiltonian and Grassmann versions of the quartic interaction differ by these $\tfrac{1}{2}$ factors, see [39, Sect.5.1], compare in particular [39, eq.(5.2)] and [39, eq.(5.9)]).
II.2 Correlation functions and Kubo conductivity
The central object of our interest is the Kubo conductivity, which is written in terms of the two-point current correlation function. We recall that the lattice current operator is defined by first promoting the Hamiltonian $\mathcal{H}_{L}$ to be $\vec{A}$-dependent, where $\vec{A}$ is an external $U(1)$ vector potential, via a replacement of the hopping parameters to $\vec{A}$-dependent hoppings via the Peierls substitution; and then by taking $\vec{\mathcal{J}}_{\vec{p}}=\delta H(\vec{A})/\delta\vec{A}_{\vec{p}}\big|_{\vec{A}=\vec{0}}$, see [32, Appendix A]; or, equivalently, it can be defined by imposing the lattice continuity equation, see [33, Sect.2.2]. The result is
$\vec{\mathcal{J}}_{\vec{p}}=\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}\hat{\psi}^{\dagger}_{\vec{k}+\vec{p}}\vec{\Gamma}(\vec{k},\vec{p})\hat{\psi}_{\vec{k}},$ (4)
where $\vec{p}\in\mathcal{D}_{L}:=\{\vec{k}=\tfrac{n_{1}}{L}\vec{G}_{1}+\tfrac{n_{2}}{L}\vec{G}_{2}:n_{i}\in\mathds{Z}\}$, and the components of the vector $\vec{\Gamma}(\vec{k},\vec{p})$ are $2\times 2$ matrices, called the bare vertex functions, such that
$\begin{split}&\vec{\Gamma}_{1,1}(\vec{k},\vec{p})=-it_{2}\sum_{j=1}^{3}\sum_{\alpha=\pm}\alpha\vec{\gamma}_{j}\eta_{\alpha\vec{\gamma}_{j}\cdot\vec{p}}e^{i\alpha(\phi-\vec{k}\cdot\vec{\gamma}_{j})}\\ &\vec{\Gamma}_{1,2}(\vec{k},\vec{p})=-it_{1}\sum_{j=1}^{3}\vec{\delta}_{j}\eta_{\vec{\delta}_{j}\cdot\vec{p}}e^{-i\vec{k}\cdot(\vec{\delta}_{j}-\vec{\delta}_{1})},\end{split}$ (5)
with $\eta_{x}:=(e^{-ix}-1)/(-ix)$, while $\vec{\Gamma}_{2,1}(\vec{k},\vec{p})=-\vec{\Gamma}_{1,2}(-\vec{k}-\vec{p},\vec{p})$, and $\vec{\Gamma}_{2,2}(\vec{k},\vec{p})=-e^{-i\vec{p}\cdot\vec{\delta}_{1}}\vec{\Gamma}_{1,1}(-\vec{k},-\vec{p})$.
In particular, the current operator at $\vec{p}=\vec{0}$ can be written as (cf. [39, Appendix A])
$\vec{\mathcal{J}}_{\vec{0}}=-\frac{1}{L^{2}}\sum_{\vec{k}\in\mathcal{B}_{L}}\tilde{\psi}^{\dagger}_{\vec{k}}\nabla_{\vec{k}}\tilde{H}^{0}(\vec{k})\tilde{\psi}_{\vec{k}},$ (6)
where $\tilde{\psi}_{\vec{k}}=\Gamma^{\dagger}_{0}(\vec{k})\hat{\psi}_{\vec{k}}$ and $\tilde{H}^{0}(\vec{k})=\Gamma^{\dagger}_{0}(\vec{k})\hat{H}^{0}(\vec{k})\Gamma_{0}(\vec{k})$. Here, $\Gamma_{0}(\vec{k})=\begin{pmatrix}1&0\\ 0&e^{-ik_{1}}\end{pmatrix}$ and $\hat{H}^{0}(\vec{k})$ is the Bloch Hamiltonian of the non-interacting system (cf. [38] and [39, Appendix B]):
$\hat{H}^{0}(\vec{k})=$ (7)
$\left(\begin{array}[]{cc}-2t_{2}\alpha_{1}(\vec{k})\cos\phi+m(\vec{k})&-t_{1}\Omega^{*}(\vec{k})\\ -t_{1}\Omega(\vec{k})&-2t_{2}\alpha_{1}(\vec{k})\cos\phi-m(\vec{k})\\ \end{array}\right),$ (10)
where
$\alpha_{1}(\vec{k})$ $=\sum_{j=1}^{3}\cos(\vec{k}\cdot\vec{\gamma}_{j}),\hskip 14.22636pt$ $m(\vec{k})=W-2t_{2}\alpha_{2}(\vec{k})\sin\phi,$
$\alpha_{2}(\vec{k})$ $=\sum_{j=1}^{3}\sin(\vec{k}\cdot\vec{\gamma}_{j}),\hskip 14.22636pt$ $\Omega(\vec{k})=1+e^{-i\vec{k}\cdot\vec{\ell}_{1}}+e^{-i\vec{k}\cdot\vec{\ell}_{2}}.$
In the thermodynamic and zero-temperature limits, after a Wick rotation of the time variable, the Kubo conductivity matrix of elements $\sigma_{ij}$ can be expressed in terms of Euclidean current-current correlations, see [39, Theorem 3.1] for a proof of the validity of the Wick rotation for the conductivity in the off-critical case, and [41, Sec.5] for a proof in the critical case. One finds that
$\sigma_{ij}=\frac{1}{|\vec{\ell}_{1}\times\vec{\ell}_{2}|}\partial^{-}\hat{K}_{ij}(0),$ (11)
where $i,j\in\{1,2\}$, $|\vec{\ell}_{1}\times\vec{\ell}_{2}|=3\sqrt{3}/2$ is the unit cell area, $\partial^{-}\hat{K}_{ij}(0)=\lim_{p_{0}\to 0^{-}}\tfrac{\hat{K}_{ij}(p_{0})-\hat{K}_{ij}(0)}{p_{0}}$ is the left derivative at $p_{0}=0$, and
$\begin{split}&\hat{K}_{ij}(p_{0})=\lim_{\beta\to\infty}\lim_{L\to\infty}\frac{1}{\beta L^{2}}\int_{0}^{\beta}dt_{1}\int_{0}^{\beta}dt_{2}\,e^{-ip_{0,\beta}(t_{1}-t_{2})}\cdot\\ &\quad\cdot\Big[\langle T\big(\mathcal{J}_{\vec{0},i}(t_{1})\mathcal{J}_{\vec{0},j}(t_{2})\big)\rangle_{\beta,L}-\langle\mathcal{J}_{\vec{0},i}\rangle_{\beta,L}\langle\mathcal{J}_{-\vec{0},j}\rangle_{\beta,L}\Big]\end{split}$ (12)
is the Euclidean current-current correlation at Matsubara frequency $p_{0}$ in the thermodynamic and zero-temperature limits. In (12), the expectation $\langle\cdot\rangle_{\beta,L}$ is computed with respect to the Gibbs measure at inverse temperature $\beta$,
$\langle\cdot\rangle_{\beta,L}=\frac{\mbox{Tr}\left\{e^{-\beta\mathcal{H}_{L}}\cdot\right\}}{\mbox{Tr}\left\{e^{-\beta\mathcal{H}_{L}}\right\}},$
and, for any $0\leq t<\beta$, $\mathcal{J}_{\vec{p},i}(t)=e^{t\mathcal{H}_{L}}\mathcal{J}_{\vec{p},i}e^{-t\mathcal{H}_{L}}$ is the imaginary time evolution of the current; moreover, $p_{0,\beta}=\tfrac{2\pi}{\beta}\lfloor\tfrac{\beta p_{0}}{2\pi}\rfloor$ is an integer multiple of $2\pi/\beta$ tending to $p_{0}$ as $\beta\to\infty$; finally, the operator $T$ inside the expectation on the RHS is the time-ordering operator, which reorders the product of the two time-dependent operators in its argument in decreasing time order.
For later reference, we also introduce the two-point Schwinger function in the thermodynamic and zero-temperature limits (i.e., the Euclidean Green’s function)
$\begin{split}\hat{S}(k_{0},\vec{k})&=\lim_{\beta\to\infty}\lim_{L\to\infty}\tfrac{1}{\beta}\int_{0}^{\beta}dt_{1}\int_{0}^{\beta}dt_{2}\sum_{\vec{x}\in\Lambda_{L}}\cdot\\ &\cdot e^{ik_{0,\beta}(t_{1}-t_{2})+i\vec{k}_{L}\cdot\vec{x}}\langle T\big(\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big)\rangle_{\beta,L},\end{split}$ (13)
where $k_{0,\beta}=\tfrac{2\pi}{\beta}(\lfloor\tfrac{\beta\omega}{2\pi}\rfloor+\tfrac{1}{2})$, $(\vec{k}_{L})_{i}=\tfrac{2\pi}{L}\lfloor\tfrac{Lk_{i}}{2\pi}\rfloor$, and, again, $\psi^{(\dagger)}_{\vec{x}}(t)=e^{t\mathcal{H}_{L}}\psi^{(\dagger)}_{\vec{x}}e^{-t\mathcal{H}_{L}}$ indicates the time evolution of the creation or annihilation fermionic operators, while $T$ is the fermionic time ordering operator such that
$T\big(\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big)=\begin{cases}\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})&\text{if $t_{1}>t_{2}$}\\ -\psi^{\dagger}_{\vec{0}}(t_{2})\psi_{\vec{x}}(t_{1})&\text{if $t_{1}\leq t_{2}$}.\end{cases}$
II.3 The non-interacting theory
Before proceeding further, it is convenient to briefly recall a few properties of the system at $U=0$, in which case the model is exactly solvable. The energy bands, i.e., the $\vec{k}$-dependent eigenvalues of the Bloch Hamiltonian (7), are
$\varepsilon_{\pm}(\vec{k})=-2t_{2}\alpha_{1}(\vec{k})\cos\phi\pm\sqrt{m(\vec{k})^{2}+t_{1}^{2}|\Omega(\vec{k})|^{2}}.$
Under the assumption $t_{2}/t_{1}<1/3$, the two bands do not overlap, and can only touch at the two Fermi points, where $\Omega(\vec{k})$ vanishes: $\vec{k}_{F}^{\,\pm}=\frac{2\pi}{3}\big(1,\pm\tfrac{1}{\sqrt{3}}\big)$. Denoting $m_{\omega}:=m(\vec{k}_{F}^{\omega})$, with $\omega\in\{+,-\}$, the two bands touch at $\vec{k}_{F}^{\omega}$ iff
$m_{\omega}(\phi,W)=W+3\sqrt{3}\omega t_{2}\sin\phi=0.$ (14)
This equation defines two curves in the $(\phi,W)$-plane, which we call the critical curves of the non-interacting theory. We fix the chemical potential at $\mu=-2t_{2}\alpha_{1}(\vec{k}_{F}^{+})\cos\phi=3t_{2}\cos\phi$, so that, if $(\phi,W)$ is on the critical curves, the system is a semimetal, while, in the complement of the critical curves, it is in an insulating (gapped) phase. In that case, the computation of the conductivity matrix leads to the values indicated in Fig. 1: the computation in the off-critical, insulating case was discussed in [39, Appendix B]; the computation of the longitudinal conductivity on the critical lines was discussed in [29, Sect.IV] (at the graphene point) and in [32, Sect.IV.B] (away from the graphene point); the computation of the transverse conductivity on the criticial lines, away from the graphene points where it is trivially zero, is discussed in Section III.1 below.
For completeness, let us conclude this section by describing the form that the two-point Schwinger function (13) takes in the non-interacting case,
$\hat{S}^{0}(k_{0},\vec{k}):=\hat{S}(k_{0},\vec{k})\big|_{U=0}=\big(-ik_{0}+\hat{H}^{0}(\vec{k})-\mu\big)^{-1},$ (15)
with $\mu=3t_{2}\cos\phi$. Its Fourier dual is denoted by $S^{0}(t_{1}-t_{2},\vec{x})=\lim_{\beta\to\infty}\lim_{L\to\infty}\langle T\big(\psi_{\vec{x}}(t_{1})\psi^{\dagger}_{\vec{0}}(t_{2})\big)\rangle_{\beta,L}\big|_{U=0}$, and its partial Fourier dual (with respect to the Matsubara frequency) by $\widetilde{S}^{0}(t,\vec{k}):=\sum_{\vec{x}\in\Lambda}e^{i\vec{k}\cdot\vec{x}}S^{0}(t,\vec{x})$, which reads
$\tilde{S}^{0}(t,\vec{k})=e^{-t(\hat{H}^{0}(\vec{k})-\mu)}\left(\mathds{1}_{\{t>0\}}P_{+}(\vec{k})-\mathds{1}_{\{t\leq 0\}}P_{-}(\vec{k})\right),$ (16)
where $P_{+}(\vec{k})$ and $P_{-}(\vec{k})$ are the projections over the upper and lower bands, respectively.
If both $m_{+}$ and $m_{-}$ are different from zero, then $\hat{S}^{0}(\mathbf{k})$ is real analytic in $\mathbf{k}\equiv(k_{0},\vec{k})$ over $\mathbb{R}\times\mathcal{B}$, and its Fourier dual $S^{0}(\mathbf{x})$ decays exponentially over $\mathbb{R}\times\Lambda$. Conversely, on the critical lines defined by (14), $\hat{S}^{0}(\mathbf{k})$ in (15) has a simple pole at $\mathbf{k}_{F}^{\omega}\equiv(0,\vec{k}_{F}^{\omega})$:
$\begin{split}&\hat{S}^{0}\left(\mathbf{k}_{F}^{\omega}+\mathbf{k}^{\prime}\right)=\\ &\left(\begin{array}[]{cc}-ik_{0}^{\prime}&\frac{3}{2}t_{1}(ik_{1}^{\prime}-\omega k_{2}^{\prime})\\ -\frac{3}{2}t_{1}(ik_{1}^{\prime}+\omega k_{2}^{\prime})&-ik_{0}^{\prime}\\ \end{array}\right)^{-1}\left(1+O(|\mathbf{k}^{\prime}|)\right),\end{split}$ (17)
as $|\mathbf{k}^{\prime}|\to 0$, and, correspondingly, $S^{0}(\mathbf{x})$ decays algebraically at large time-space distances, asymptotically bounded from above and below by a constant times $|\mathbf{x}|^{-2}$.
II.4 Main results: the Hall conductivity in the critical regime
As anticipated in the introduction, our main result concerns the transverse conductivity $\sigma_{12}=-\sigma_{21}$, and can be stated as follows.
Theorem II.1.
There exists $U_{0}>0$, independent of $W,\phi$, such that, for any $|U|<U_{0}$, there exist two functions $\mathfrak{d}(U,W,\phi),\mathfrak{z}(U,W,\phi)$, analytic in $U$, vanishing at $U=0$, and continuously differentiable in $W,\phi$, such that, if the chemical potential $\mu$ is fixed at the value $3t_{2}\cos\phi-\mathfrak{z}(U,W,\phi)$, the Hall conductivity of the interacting Haldane model (2) reads:
$\sigma_{12}=\tfrac{1}{4\pi}\Big[\textup{ sgn}\left(m^{R}_{+}(U,W,\phi)\right)-\textup{ sgn}\left(m^{R}_{-}(U,W,\phi)\right)\Big],$ (18)
where $m^{R}_{\pm}(U,W,\phi)\equiv W\pm 3\sqrt{3}t_{2}\sin\phi\pm\mathfrak{d}(U,W,\phi)$, with the understanding that $\textup{sgn}(0)=0$.
As discussed in [32, 33], the functions $m^{R}_{\pm}$ have the meaning of dressed, renormalized, masses. The system is critical, i.e., its Euclidean correlation functions decay polynomially to zero at large space-time separation, iff either of the two masses vanishes.
The proof of the theorem is constructive, that is, it is based on an algorithm allowing one to compute $U_{0}$, as well as the Taylor coefficients of the analytic functions $\mathfrak{d}$ and $\mathfrak{z}$, representing the interaction-induced shifts of the mass and of the chemical potential, respectively (see the discussion after eq.(28), and [33, Lemma 4.2] for additional details; see also [32, Section III.E] for the explicit lowest-order computation of $\mathfrak{z}$ in the case of ultra-local Hubbard interaction). The value of $U_{0}$ provides an estimate on the strength of the interaction beyond which new physics appears. However, in this paper we do not attempt to evaluate it explicitly, because we do not expect that our proof can provide a realistic value for the transition strength. A numerically more refined scheme and a computer-assisted proof would be required for this purpose, but this goes beyond the scope of this work.
The proof of Theorem II.1 in the off-critical case, when both $m^{R}_{+}(U,W,\phi)$ and $m^{R}_{-}(U,W,\phi)$ are different from zero, was treated in [33].
The new case proved in this paper is the critical case where either $m^{R}_{+}(U,W,\phi)$ or $m^{R}_{-}(U,W,\phi)$ vanishes (the case where both simultaneously vanish is easier, and was treated in [29, 32]): this was precisely the missing case needed to complete the picture of the conductivity matrix displayed in Fig. 2. The reason why the computation of the critical $\sigma_{12}$ remained an open problem is that none of the methods used in the earlier works [32, 33] is directly applicable in the present case.
In [33], we proved the quantization of $\sigma_{12}$ in the off-critical case via a combination of Ward identities and Schwinger–Dyson equations, which implies the vanishing of the non-universal corrections to the conductivity provided that the current-current correlations in momentum space are three times differentiable, see [33, Sect.3.1].
Such a differentiability condition fails on the critical line, where the derivative of the Euclidean current-current correlation, $\tfrac{\partial}{\partial p_{0}}\hat{K}_{ij}(p_{0})$, is dimensionally logarithmically divergent as $p_{0}\to 0$ (note that $\sigma_{ij}$ is proportional precisely to the right derivative at $p_{0}=0$ of such an a priori singular quantity). Therefore, at criticality, one needs to proceed in a different way. An effective strategy, which works well in the case of the longitudinal conductivity [29, 32], is to rewrite $\hat{K}_{ij}(p_{0})$ as the sum of a singular contribution, coming from the lowest-order Feynman diagram (the ‘polarization bubble’) with dressed, renormalized, vertex functions, and of a regular one, given by the convergent sum of all the dressed diagrams with at least one interaction vertex; note, in fact, that the quartic interaction is dimensionally irrelevant in the infrared, which induces a dimensional gain with respect to the naive power counting on all the interaction corrections beyond the dressed polarization bubble; this, in turn, implies that the such higher order interaction corrections sum up to a quantity that is continuously differentiable at $p_{0}=0$. On the other hand, elementary parity considerations show that the longitudinal current-current correlation, $\hat{K}_{ii}(p_{0})$, is even in $p_{0}$, and so are its singular and regular parts, separately. Therefore, the derivative of the regular part of $\hat{K}_{ii}(p_{0})$ at $p_{0}=0$ is readily zero, and so is its contribution to the critical longitudinal conductivity. On the other hand, the contribution to $\partial\hat{K}_{ii}(p_{0})$ as $p_{0}\to 0^{+}$, coming from the polarization bubble can be evaluated explicitly and gives $1/8$ per Dirac cone, see [32, Sect.IV.B].
Unfortunately, a direct application of this strategy fails for the critical transverse conductivity. In fact, by parity reasons, it turns out that things go the other way round for the off-diagonal components of the conductivity matrix: the singular contribution from the polarization bubble vanishes and the whole contribution to $\sigma_{12}$ comes from the regular part. However, this regular part is not explicit: it is the convergent sum of infinitely many Feynman diagrams and there is no hope to compute the sum directly.
In this paper, we show how to compute the critical transverse conductivity via two different strategies, different from those of earlier papers. The first one, discussed in Section III.1, applies to the non-interacting theory, and is based on methods closely resembling the ones used in the off-critical case (see e.g. [39, Appendix B]), which admit an explicit interpretation in terms of the Berry curvature of the Bloch bundle. The second proof, discussed in Sections III.2 and IV), is more general, valid both in the non-interacting and interacting cases. It consists in comparing the critical transverse conductivity with an appropriate average of the off-critical values of $\sigma_{12}$ inside and outside the critical curve, at a distance $\epsilon$ to be eventually sent to zero. The difference between the critical $\sigma_{12}$ and the average of its off-critical values at distance $\epsilon$ can also be decomposed in its singular and regular parts via a multiscale renormalization group computation, which can both be shown to vanish as the regularization parameter $\epsilon$ goes to zero, either via a direct computation (as far as the singular contribution from the polarization bubble is concerned) or via dimensional bounds (as far as the regular part from the higher order interaction corrections is concerned).
III The critical transverse conductivity
III.1 The non-interacting case
In this section we compute $\sigma_{12}$ in the critical, non-interacting case, and prove in this setting the validity of (18). We let $U=0$ and, without loss of generality, we focus on the critical line $m_{-}=0$, with $\phi\in(0,\pi)$: the cases where $m_{+}=0$ and/or $\phi\in(-\pi,0)$ are equivalent thanks to the symmetries of the model, see [32, Sect.III.B].
Our starting point is Kubo’s formula (11). Plugging (6) into (12) and using Wick’s rule, we find
$\begin{split}&\hat{K}_{ij}(p_{0})=-\int_{\mathds{R}}dt\int_{\mathcal{B}}\frac{d\vec{k}}{|\mathcal{B}|}e^{-ip_{0}t}\\ &\cdot\mbox{Tr}\left\{\tilde{g}(-t,\vec{k})\partial_{k_{i}}\tilde{H}^{0}(\vec{k})\tilde{g}(t,\vec{k})\partial_{k_{j}}\tilde{H}^{0}(\vec{k})\right\}.\end{split}$ (19)
Here,
$\tilde{g}(t,\vec{k}):=e^{-t\tilde{H}^{0}(\vec{k})}\left(\mathds{1}_{\{t>0\}}\tilde{P}_{+}(\vec{k})-\mathds{1}_{\{t\leq 0\}}\tilde{P}_{-}(\vec{k})\right),$
and $\tilde{P}_{\pm}(\vec{k})=\Gamma^{\dagger}_{0}(\vec{k})P_{\pm}(\vec{k})\Gamma_{0}(\vec{k})$, with $P_{\pm}(\vec{k})$ the projections on the upper and lower bands, respectively. Plugging (19) into (11) we find that
$\begin{split}&\sigma_{ij}=-\lim_{p_{0}\rightarrow 0^{-}}\int_{\mathbb{R}}dt\int_{\mathcal{B}}\frac{d\vec{k}}{(2\pi)^{2}}\left(\frac{e^{-ip_{0}t}-1}{p_{0}}\right)\\ &\cdot\mbox{Tr}\left\{\tilde{g}(-t,\vec{k})\partial_{k_{i}}\tilde{H}^{0}(\vec{k})\tilde{g}(t,\vec{k})\partial_{k_{j}}\tilde{H}^{0}(\vec{k})\right\},\end{split}$ (20)
where we used the fact that $|\vec{\ell}_{1}\times\vec{\ell}_{2}||\mathcal{B}|=(2\pi)^{2}$. By separating the integral over $(0,+\infty)$ from the one over $(-\infty,0)$ in (20) and by changing variable $t\rightarrow-t$ in the latter, we obtain
$\begin{split}&\sigma_{ij}=\lim_{p_{0}\rightarrow 0^{-}}\int_{0}^{\infty}dt\int_{\mathcal{B}}\frac{d\vec{k}}{(2\pi)^{2}}\\ &\cdot\left[\left(\frac{e^{-ip_{0}t}-1}{p_{0}}\right)f_{ij}(t,\vec{k})+\left(\frac{e^{ip_{0}t}-1}{p_{0}}\right)f_{ji}(t,\vec{k})\right],\end{split}$ (21)
with
$\begin{split}&f_{ij}(t,\vec{k})\\ &:=\mbox{Tr}\Big\{e^{t\tilde{H}^{0}(\vec{k})}\tilde{P}_{-}(\vec{k})\partial_{k_{i}}\tilde{H}^{0}(\vec{k})e^{-t\tilde{H}^{0}(\vec{k})}\tilde{P}_{+}(\vec{k})\partial_{k_{j}}\tilde{H}^{0}(\vec{k})\Big\}.\end{split}$ (22)
We are studying the critical line $m_{-}=0$ and $\phi\in(0,\pi)$, which means that the energy gap only closes at the Fermi point $\vec{k}_{F}^{-}$. It is therefore convenient to split the integral over momenta in (21) into an integral over a small ball centered at $\vec{k}_{F}^{-}$, and an integral over its complement. That is, we introduce $B_{\epsilon}:=\left\{\vec{k}\in\mathcal{B}:|\vec{k}-\vec{k}_{F}^{-}|\leq\epsilon\right\}$ with $\epsilon\ll 1$, and write
$\sigma_{12}=\sigma_{12}^{(a)}+\sigma_{12}^{(b)},$
where $\sigma_{12}^{(a)}$ and $\sigma_{12}^{(b)}$ are given by the RHS of (21), with the integral over $d\vec{k}$ restricted to $B_{\epsilon}$ and $\mathcal{B}\setminus B_{\epsilon}$ respectively.
The contribution $\sigma_{12}^{(a)}$. We focus on the first term on the RHS of (21). Using (22) and recalling the definition of $\Gamma_{0}(\vec{k})$ (see the lines after (6)), we rewrite
$\begin{split}&f_{12}(t,\vec{k})=e^{-2t\sqrt{m(\vec{k})^{2}+t_{1}^{2}|\Omega(\vec{k})|^{2}}}\\ &\cdot\left(\mbox{Tr}\left\{P_{-}(\vec{k})\partial_{k_{1}}\hat{H}^{0}(\vec{k})P_{+}(\vec{k})\partial_{k_{2}}\hat{H}^{0}(\vec{k})\right\}\right.\\ &\left.\ \ \ +\mbox{Tr}\left\{P_{-}(\vec{k})[A_{1},\hat{H}^{0}(\vec{k})]P_{+}(\vec{k})\partial_{k_{2}}\hat{H}^{0}(\vec{k})\right\}\right),\end{split}$ (23)
with $A_{1}=\Gamma_{0}(\vec{k})\partial_{k_{1}}\Gamma^{\dagger}_{0}(\vec{k})=\left(\begin{array}[]{cc}0&0\\ 0&i\end{array}\right)$.
We expand the following quantities around $\vec{k}_{F}^{-}$,
$\begin{split}&\hat{H}^{0}(\vec{k}_{F}^{-}+\vec{k}^{\prime})-\mu=-\frac{3}{2}t_{1}\sigma_{2}k_{1}^{\prime}+\frac{3}{2}t_{1}\sigma_{1}k_{2}^{\prime}+O(|\vec{k}^{\prime}|^{2})\\ &P_{-}(\vec{k}_{F}^{-}+\vec{k}^{\prime})=\frac{1}{2}\left(\begin{array}[]{cc}1&-ie^{-i\text{arg}(\vec{k}^{\prime})}\\ ie^{i\text{arg}(\vec{k}^{\prime})}&1\end{array}\right)+O(|\vec{k}^{\prime}|)\\ &P_{+}(\vec{k}_{F}^{-}+\vec{k}^{\prime})=\frac{1}{2}\left(\begin{array}[]{cc}1&ie^{-i\text{arg}(\vec{k}^{\prime})}\\ -ie^{i\text{arg}(\vec{k}^{\prime})}&1\end{array}\right)+O(|\vec{k}^{\prime}|)\\ &\sqrt{m(\vec{k}_{F}^{-}+\vec{k}^{\prime})^{2}+t_{1}^{2}|\Omega(\vec{k}_{F}^{-}+\vec{k}^{\prime})|^{2}}=\frac{3}{2}t_{1}|\vec{k}^{\prime}|+O(|\vec{k}^{\prime}|^{2}),\end{split}$
where $\sigma_{1},\sigma_{2},\sigma_{3}$ are the Pauli matrices and $\arg(\vec{k}^{\prime})$ is the argument of $k_{1}^{\prime}+ik_{2}^{\prime}$. This allows us to expand (23) and estimate
$\begin{split}&\left|f_{12}(t,\vec{k})+\frac{9}{8}t_{1}^{2}\sin(2\text{arg}(\vec{k}^{\prime}))e^{-3tt_{1}|\vec{k}^{\prime}|}\right|\\ &\leq O\left(|\vec{k}^{\prime}|e^{-tt_{1}|\vec{k}^{\prime}|}\right)+O\left(t|\vec{k}^{\prime}|^{2}e^{-tt_{1}|\vec{k}^{\prime}|}\right),\end{split}$ (24)
for $|\vec{k}^{\prime}|$ sufficiently small. A similar computation for $f_{21}(t,\vec{k})$ gives the same leading order, so that
$\begin{split}\sigma_{12}^{(a)}&=O(\epsilon)-\lim_{p_{0}\rightarrow 0^{-}}\int_{0}^{\infty}dt\int_{B_{\epsilon}(\vec{0})}\frac{d\vec{k}^{\prime}}{(2\pi)^{2}}\\ &\cdot\left(\frac{e^{ip_{0}t}+e^{-ip_{0}t}-2}{p_{0}}\right)\frac{9}{8}t_{1}^{2}\sin(2\text{arg}(\vec{k}^{\prime}))e^{-3tt_{1}|\vec{k}^{\prime}|},\end{split}$ (25)
where the term bounded as $O(\epsilon)$ was obtained from the RHS of (24). Using polar coordinates $(|\vec{k}^{\prime}|,\text{arg}(\vec{k}^{\prime}))$, we see that the integral in (25) vanishes by parity, and so $\sigma_{12}^{(a)}=O(\epsilon)$.
The contribution $\sigma_{12}^{(b)}$. We closely follow Appendices A and B of [39], which discuss the computation of $\sigma_{12}$ in the non-critical case. The reason these methods apply in the present case is that the energy bands do not touch on $\mathcal{B}\setminus B_{\epsilon}$ (recall that $\sigma_{12}^{(b)}$ is the RHS of (21) with the integration of $\vec{k}$ restricted to $\mathcal{B}\setminus B_{\epsilon}$).
Note that for $\epsilon>0$ fixed, $f_{12}(t,\vec{k})$ and $f_{21}(t,\vec{k})$ decay exponentially in $t$, uniformly in $\vec{k}\in\mathcal{B}\setminus B_{\epsilon}$. Hence, we can bring the limit inside the integral over $t$,
$\sigma_{12}^{(b)}=-i\int_{0}^{\infty}dt\ t\int_{\mathcal{B}\setminus B_{\epsilon}}\frac{d\vec{k}}{(2\pi)^{2}}\left[f_{12}(t,\vec{k})-f_{21}(t,\vec{k})\right].$ (26)
We also have that (cf. appendix A of [39])
$\begin{split}&f_{ij}(t,\vec{k})\\ &=\partial^{2}_{t}\mbox{Tr}\left\{e^{t\tilde{H}^{0}(\vec{k})}\tilde{P}_{-}(\vec{k})\partial_{k_{i}}\tilde{P}_{-}(\vec{k})e^{-t\tilde{H}^{0}(\vec{k})}\partial_{k_{j}}\tilde{P}_{-}(\vec{k})\right\},\end{split}$
where the trace decays exponentially to zero as $t\to\infty$, uniformly in $\vec{k}$. Using partial integration, this implies
$\int_{0}^{\infty}dt\ tf_{ij}(t,\vec{k})=\mbox{Tr}\left\{\tilde{P}_{-}(\vec{k})\partial_{k_{i}}\tilde{P}_{-}(\vec{k})\partial_{k_{j}}\tilde{P}_{-}(\vec{k})\right\},$
and using again the definition of $\Gamma_{0}(\vec{k})$, we see that (26) becomes
$\begin{split}&\sigma_{12}^{(b)}=-i\int_{\mathcal{B}\setminus B_{\epsilon}}\frac{d\vec{k}}{(2\pi)^{2}}\\ &\left(\mbox{Tr}\left\{P_{-}(\vec{k})[\partial_{k_{1}}P_{-}(\vec{k}),\partial_{k_{2}}P_{-}(\vec{k})]\right\}-\partial_{k_{2}}\mbox{Tr}\left\{A_{1}P_{-}(\vec{k})]\right\}\right),\end{split}$ (27)
with $A_{1}$ as defined below (23). We now write $P_{-}(\vec{k})$ as $v_{-}(\vec{k})v^{\dagger}_{-}(\vec{k})$, with
$v_{-}(\vec{k})=\frac{1}{N(\vec{k})}\left(\begin{array}[]{cc}t_{1}\Omega^{*}(\vec{k})\\ \sqrt{m(\vec{k})^{2}+t_{1}^{2}|\Omega(\vec{k})|^{2}}+m(\vec{k})\end{array}\right),$
and $N(\vec{k})$ a normalization factor, equal to
$\left[2m(\vec{k})^{2}+2t_{1}^{2}|\Omega(\vec{k})|^{2}+2m(\vec{k})\sqrt{m(\vec{k})^{2}+t_{1}^{2}|\Omega(\vec{k})|^{2}}\right]^{\frac{1}{2}}.$
It is easily checked that $v_{-}$ is real analytic over the whole Brillouin zone, except at the point $\vec{k}_{F}^{-}$. Note that
$\mbox{Tr}\big\{P_{-}(\vec{k})[\partial_{k_{1}}P_{-}(\vec{k}),\partial_{k_{2}}P_{-}(\vec{k})]\big\}=\nabla\times\langle v_{-}(\vec{k}),\nabla v_{-}(\vec{k})\rangle,$
which is proportional to the Berry curvature. By Stokes’ theorem, (27) becomes
$\begin{split}&\sigma_{12}^{(b)}\\ &=\frac{i}{(2\pi)^{2}}\int_{\partial B_{\epsilon}}d\vec{k}\cdot\left[\langle v_{-}(\vec{k}),\nabla v_{-}(\vec{k})\rangle+\mbox{Tr}\big\{A_{1}P_{-}(\vec{k})\big\}\left(\begin{smallmatrix}1\\ 0\\ \end{smallmatrix}\right)\right]\\ &=\frac{i}{(2\pi)^{2}}\int_{\partial B_{\epsilon}}d\vec{k}\cdot\langle v_{-}(\vec{k}),\nabla v_{-}(\vec{k})\rangle+O(\epsilon),\end{split}$
where the line integral over the boundary of $B_{\epsilon}$ is run counterclockwise, and we used that $A_{1}P_{-}(\vec{k})$ is uniformly bounded in matrix norm for the second step. Setting $\vec{k}-\vec{k}_{F}^{-}=\epsilon(\cos\vartheta,\sin\vartheta)$, we can expand
$v_{-}(\vec{k})=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}-ie^{-i\vartheta}\\ 1\end{array}\right)+O(\epsilon).$
Combining the previous two equations, it follows that $\sigma_{12}^{(b)}=1/4\pi+O(\epsilon)$, so that by letting $\epsilon\rightarrow 0$ we obtain the desired result
$\sigma_{12}=\sigma_{12}^{(a)}+\sigma_{12}^{(b)}=\frac{1}{4\pi}.$
III.2 The general, interacting, case
The computation of $\sigma_{12}$ on the critical line, presented in the previous section, is based on an explicit evaluation of the current-current response function, which cannot straightforwardly be generalized to the interacting model. In that case, correlation functions cannot be written in closed form; rather, they can be expressed in the form of infinite, yet convergent, series expansions, as reviewed below. In order to compute $\sigma_{12}$ on the dressed critical line, we use the fact that $\sigma_{12}$ is quantized and universal over each of the four disconnected regions of the interacting phase diagram shown in Fig.2, as proved in [33]. As anticipated above, our goal is to show that the difference between the critical transverse conductivity and the symmetric average of the off-critical one, at a distance $\epsilon$ inside or outside the critical curve, vanishes as $\epsilon\to 0$, which implies the new part of our main result.
Let us describe our strategy more precisely. As before, without loss of generality, we focus on the critical line defined by $m^{R}_{-}(U,W,\phi)=0$ with $\phi\in(0,\pi)$. As in [33, Sect.2.3.1], we introduce the reference Hamiltonian, defined in terms of $U,m_{-}^{R},\phi$ as follows:
$\begin{split}&\mathcal{H}^{R}_{L}(U,m^{R}_{-},\phi):=\\ &\mathcal{H}_{0,L}\left(m^{R}_{-}+3\sqrt{3}t_{2}\sin\phi,\phi\right)-3t_{2}\cos\phi\ \mathcal{N}_{L}+U\mathcal{V}_{L}\\ &\ \ \ +\xi(U,m^{R}_{-},\phi)\sum_{\vec{x}\in\Lambda_{L}}\psi^{\dagger}_{\vec{x}}\psi_{\vec{x}}+\delta(U,m^{R}_{-},\phi)\sum_{\vec{x}\in\Lambda_{L}}\psi^{\dagger}_{\vec{x}}\sigma_{3}\psi_{\vec{x}}.\end{split}$ (28)
where $\xi$ and $\delta$ are the chemical potential and mass counterterms, respectively, and $\sigma_{3}$ is the third Pauli matrix. The functions $\xi$ and $\delta$ are analytic in $U$, and should be thought of, respectively, as the shifts in $\mu$ and $W$ caused by the interaction (and thus vanishing for $U=0$). They are fixed so that the RG flow of the theory approaches a Gaussian fixed point in the infrared, characterized by a renormalized mass equal to $m^{R}_{-}$ (cf. [33, Lemma 4.2]), and by a semimetallic behavior at $m^{R}_{-}=0$. Note that $\mathcal{H}^{R}_{L}(U,m^{R}_{-},\phi)$ is a perturbation of the non-interacting Hamiltonian $\mathcal{H}_{0,L}\left(m^{R}_{-}+3\sqrt{3}t_{2}\sin\phi,\phi\right)-3t_{2}\cos\phi\ \mathcal{N}_{L}$, which is critical on the ‘correct’, interacting critical line $m^{R}_{-}=0$. We require that the reference Hamiltonian coincides with $\mathcal{H}_{L}$ at parameters $W,\phi,\mu,U$, so that
$\begin{split}W&=m^{R}_{-}+3\sqrt{3}t_{2}\sin\phi+\delta(U,m^{R}_{-},\phi)\\ \mu&=3t_{2}\cos\phi-\xi(U,m^{R}_{-},\phi).\end{split}$ (29)
The function $m^{R}_{-}(U,W,\phi)$ that appears in Theorem II.1 is nothing but the solution of the equation given by the first of (29), and the functions $\mathfrak{z}(U,W,\phi),\mathfrak{d}(U,W,\phi)$ correspond to $\xi(U,m^{R}_{-},\phi)$ and $\delta(U,m^{R}_{-},\phi)$ respectively, computed at $m^{R}_{-}=m^{R}_{-}(U,W,\phi)$. In other words, the reference Hamiltonian is just a rewriting of the original interacting Hamiltonian, re-expressed in terms of new parameters, in particular of the dressed, renormalized mass, which is a physically more natural parameter to deal with, particularly if we intend to study the system at, or close to, the dressed critical line.
The multiscale construction of the counterterms and of the dressed critical line, via the inversion of (29), has been discussed in detail in [33], and will be reviewed below, for the purpose of deriving the new results required for the proof of the main theorem in the critical case.
By construction, the critical transverse conductivity we are interested in is equivalent to the transverse conductivity of the reference model at $m^{R}_{-}=0$, which we denote by
$\sigma^{R}_{12}(U,m^{R}_{-},\phi)\big|_{m^{R}_{-}=0}\equiv\sigma^{R}_{12}(U,0,\phi).$ (30)
Our goal is to show that this quantity is equal to $1/4\pi$, exactly as in the non-interacting case. The key ingredient used to prove this is a representation of the transverse conductivity of the reference model at, or close to, the massless case, following from the multiscale expansion reviewed below and summarized in the following proposition.
Proposition III.1.
There exists $U_{0}>0$ such that for $|U|\leq U_{0}$, $\phi\in(0,\pi)$, and any sufficiently small, positive, $\epsilon$, the transverse conductivity of the reference model $\sigma_{12}^{R}(U,m^{R}_{-},\phi)$ with $\epsilon\leq m^{R}_{-}\leq\epsilon$ can be decomposed as follows,
$\sigma_{12}^{R}(U,m^{R}_{-},\phi)=\sigma_{12}^{R,(1)}(U,m^{R}_{-},\phi)+\sigma_{12}^{R,(2)}(U,m^{R}_{-},\phi).$ (31)
Here, the first term (the ‘singular part’) is
$\sigma_{12}^{R,(1)}(U,m,\phi)=\gamma^{2}\lim_{p_{0}\to 0^{-}}\frac{1}{p_{0}}\int_{\mathcal{A}_{\epsilon}}\frac{d\mathbf{k}}{(2\pi)^{3}}\cdot$ (32)
$\qquad\qquad\cdot\textup{Tr}\left\{\sigma_{2}S_{m}(\mathbf{k})\sigma_{1}\big({S}_{m}(\mathbf{k}+\mathbf{p})-S_{m}(\mathbf{k})\big)\right\},$
where $\gamma$ is an analytic function of $U$, equal to $\tfrac{3}{2}t_{1}$ at $U=0$; $\mathcal{A}_{\epsilon}$ is the 3D ball of radius $\sqrt{\epsilon}$ centered at the origin; $\sigma_{1}$ and $\sigma_{2}$ are the first two Pauli matrices; $\mathbf{p}=(p_{0},0,0)$, and the ‘dressed propagator’ $S_{m}(\mathbf{k})$ at $\mathbf{k}=(k_{0},k_{1},k_{2})$ is
${S}_{m}(\mathbf{k})=\left(\begin{array}[]{cc}-iZ_{1}k_{0}+m&v(ik_{1}+k_{2})\\ v(-ik_{1}+k_{2})&-iZ_{2}k_{0}-m\\ \end{array}\right)^{-1},$ (33)
with $Z_{1},Z_{2},v$ three ‘dressed parameters’ that are analytic functions of $U$, equal to $1$, 11, $\frac{3}{2}t_{1}$ at $U=0$, respectively.
Moreover, the second term in (31) (the ‘regular part’) is continuous at zero, in the sense that
$\begin{split}\lim_{\epsilon\rightarrow 0^{+}}\sigma_{12}^{R,(2)}(U,\pm\epsilon,\phi)=\sigma_{12}^{R,(2)}(U,0,\phi).\end{split}$ (34)
The ‘singular’ contribution $\sigma_{12}^{R,(1)}$ is nothing but the dressed lowest order diagram contribution in renormalized perturbation theory, which is the dominant contribution at low momenta. The dressing consists in a finite (analytic!) renormalization of the parameters $Z_{1},Z_{2},v$ in the interacting propagator, see (33), and of the vertex functions in directions $1$ and $2$, equal to $\gamma\sigma_{2}$ and $\gamma\sigma_{1}$ respectively. In (32), $\mathbf{k}$ represents the quasi-momentum relative to the Fermi point $\vec{k}^{-}_{F}$; the restriction of $\mathbf{k}$ to $\mathcal{A}_{\epsilon}$ corresponds to the fact that the singular part only includes contributions from the infrared modes, at low quasi-momenta. The regular part of the conductivity includes all the contributions at larger quasi-momenta, as well as the (convergent!) sum of all the Feynman diagrams in renormalized perturbation theory involving a quartic electron-electron interaction; since the quartic interaction is irrelevant in the infrared (in an RG sense), all such terms are dimensionally better behaved in the infrared, as compared to the dominant contribution leading to (32), and this is ultimately the reason why the regular part is continuous on the critical line, as per (34).
Proposition III.1 is a corollary of the multiscale construction of the model, and it will be proved in the next section. Before we discuss it, let us prove here that the proposition readily implies the desired result for the critical conductivity. In order to compute (30), we rewrite:
$\begin{split}&\sigma_{12}^{R}(U,0,\phi)=\tfrac{1}{2}\lim_{\epsilon\to 0^{+}}\big(\sigma_{12}^{R}(U,\epsilon,\phi)+\sigma_{12}^{R}(U,-\epsilon,\phi)\big)\\ &\quad+\lim_{\epsilon\to 0^{+}}\big[\sigma_{12}^{R}(U,0,\phi)-\tfrac{1}{2}\sigma_{12}^{R}(U,\epsilon,\phi)-\tfrac{1}{2}\sigma_{12}^{R}(U,-\epsilon,\phi)\big].\end{split}$ (35)
Thanks to the validity of Theorem II.1 in the off-critical case, already proved in [33], the first term on the RHS equals $\tfrac{1}{4\pi}$ for all $\phi\in(0,\pi)$. We want to show that the second line vanishes, thanks to Proposition III.1. We rewrite the conductivity as in (31) and then, using the continuity property (34), notice that the only possible contribution to the second line of (35) comes from $\sigma_{12}^{R,(1)}$. In order to compute the contribution from $\sigma_{12}^{R,(1)}$, we recall (32), use (33), and find that
$\begin{split}&\frac{1}{p_{0}}\textup{Tr}\left\{\sigma_{2}S_{m}(\mathbf{k})\sigma_{1}\big({S}_{m}(\mathbf{k}+\mathbf{p})-S_{m}(\mathbf{k})\big)\right\}=\\ &=\frac{4v^{2}[Z_{1}Z_{2}(2k_{0}+p_{0})-im(Z_{1}-Z_{2})]k_{1}k_{2}}{D_{m}^{2}(\mathbf{k})D_{m}(\mathbf{k}+\mathbf{p})}\\ &-\frac{m(Z_{1}+Z_{2})}{D_{m}(\mathbf{k})D_{m}(\mathbf{k}+\mathbf{p})},\end{split}$ (36)
where $D_{m}(\mathbf{k})=Z_{1}Z_{2}k_{0}^{2}+v^{2}(k_{1}^{2}+k_{2}^{2})+m^{2}-im(Z_{1}-Z_{2})k_{0}$; next, we plug (36) into (32), thus getting, for $m=0$ and $m=\pm\epsilon$ respectively:
$\sigma_{12}^{R,(1)}(U,0,\phi)=0$
and
$\sigma_{12}^{R,(1)}(U,\pm\epsilon,\phi)=\mp\epsilon(Z_{1}+Z_{2})\int_{\mathcal{A}_{\epsilon}}\frac{d\mathbf{k}}{(2\pi)^{3}}\frac{1}{\big[D_{\pm\epsilon}(\mathbf{k})\big]^{2}}.$
The value of the integral in the last equation is independent of the choice of the sign in the subscript of $D_{\pm\epsilon}(\mathbf{k})$. Therefore, $\sigma_{12}^{R,(1)}(U,\epsilon,\phi)+\sigma_{12}^{R,(1)}(U,-\epsilon,\phi)=0$, which implies that the second line of (35) vanishes as announced.
Remark 1.
Note that the fact that $\sigma_{12}^{R,(1)}(U,0,\phi)=\frac{1}{2}\big(\sigma_{12}^{R,(1)}(U,\epsilon,\phi)+\sigma_{12}^{R,(1)}(U,-\epsilon,\phi)\big)=0$ ultimately follows from an emergent parity symmetry of the effective relativistic infrared theory, which implies that $\sigma_{12}^{R,(1)}(U,\epsilon,\phi)$ is odd in $\epsilon$, thus enforcing that the right average to use in (35) is the standard arithmetic one, rather than other, more general, weighted averages.
Putting things together, we proved that
$\sigma_{12}^{R}(U,0,\phi)=\frac{1}{4\pi},$
as desired. We are left with discussing the proof Proposition III.1, which is described in the next section.
IV Proof of Prop. III.1
In this section we prove Proposition III.1, whose result, as already mentioned, is a corollary of the multiscale RG construction of the interacting ground state, already discussed in detail in [32, 33]. In order to explain the origin of the decomposition in (31) and of the continuity of $\sigma_{12}^{R,(2)}$ at $m^{R}_{-}=0$, we first review the main aspects of the RG construction, at a level of detail sufficient for us to state and justify the main computations and estimates behind the proof of Proposition III.1, referring the reader to [32, 33] for additional details. Next, we use the output of the RG construction to prove the main content of the proposition of interest, that is the continuity bound (34). We assume, as before, that $\phi\in(0,\pi)$, and we take $|m_{-}^{R}|$ sufficiently small, so that $m_{+}^{R}:=m_{-}^{R}+6\sqrt{3}t_{2}\sin\phi>|m_{-}^{R}|$. We stress that the method we use not only allows us to account for the effect of interactions, that is, it is robust under smooth modifications of the quartic part of the Hamiltonian (i.e., the interaction), but it is also flexible enough to deal with smooth modifications of the quadratic part of the Hamiltonian (i.e., the hopping): as clear from the discussion below, the only thing we really use is the structure of the relevant and marginal terms, see, e.g., eq.(43) below, which ultimately follows from the symmetry properties of the model and the number of (quasi-)massless degrees of freedom at or close to the transition line. Therefore, even if the proof is spelled out only in the case of the interacting Haldane model, we expect it can be generalized to a larger class of interacting electron systems with critical semimetallic state; this will be the subject of a future publication.
IV.1 Review of the RG construction
The starting point is a reformulation of the reference model (28) in terms of an interacting Grassmann integral. The generating functional $\mathcal{W}$ of multipoint field and current correlations at inverse temperature $\beta$ and finite volume can be written as
$\mathcal{W}(\varphi,A):=\log\int P(d\psi)e^{-V(\psi)+(\psi,\varphi)+(j,A)},$ (37)
where $\psi^{\pm}_{\mathbf{x},\rho}$ with $\mathbf{x}\in[0,\beta)\times\Lambda_{L}$ and $\rho\in\{1,2\}$ are Grassmann variables, $\varphi$ is a Grassmann field conjugated to $\psi$, and $A$ is a real field conjugated to the current $j$, to be defined momentarily. In (37), $P(d\psi)$ is the Gaussian Grassmann integration with propagator
$g(\mathbf{x}-\mathbf{y})=\int\frac{d\mathbf{k}}{2\pi|\mathcal{B}|}e^{-i\mathbf{k}\cdot(\mathbf{x}-\mathbf{y})}\hat{S}^{0}(\mathbf{k}),$ (38)
where $\mathbf{x}=(x_{0},\vec{x})$, $\mathbf{k}=(k_{0},\vec{k})$, $\hat{S}^{0}(\mathbf{k})$ was defined in (15) and $\int\frac{d\mathbf{k}}{2\pi|\mathcal{B}|}$ is shorthand for the Riemann sum $\tfrac{1}{\beta L^{2}}\sum_{\mathbf{k}}$ with the sum over $\mathbf{k}$ running over $\tfrac{2\pi}{\beta}\mathbb{Z}\times\mathcal{B}_{L}$, see (1). Moreover,
$\begin{split}V(\psi)&=\int_{0}^{\beta}dx_{0}\Big[U\sum_{\begin{subarray}{c}\vec{x},\vec{y}\in\Lambda_{L}\\ \rho,\rho^{\prime}=1,2\end{subarray}}n_{(x_{0},\vec{x}),\rho}v_{\rho,\rho^{\prime}}(\vec{x}-\vec{y})n_{(x_{0},\vec{y}),\rho^{\prime}}\\ &+\sum_{\vec{x}\in\Lambda_{L}}\big(\xi(n_{\mathbf{x},1}+n_{\mathbf{x},2})+\delta(n_{\mathbf{x},1}-n_{\mathbf{x},2})\big)\Big],\end{split}$ (39)
with $n_{\mathbf{x},\rho}=\psi^{+}_{\mathbf{x},\rho}\psi^{-}_{\mathbf{x},\rho}$, and $\xi=\xi(U,m^{R}_{-},\phi)$, $\delta=\delta(U,m^{R}_{-},\phi)$ are the same as in (28). Finally,
[…]
The generating functional can be calculated in the thermodynamic and zero-temperature limits using the constructive RG approach described in [32, Sect.III.C] and in [33, Sect.4]. This allowed us to prove that the thermodynamic functions, such as the free energy per site, and the Euclidean correlation functions, also known as Schwinger functions, are analytic functions of $U$, uniformly in $\beta,L$, and that their limits as $L\to\infty$ first and then $\beta\to\infty$ are analytic in $U$ in a small complex neighborhood of the origin. Moreover, the method also allows us to compute the asymptotic behavior of correlations at large space-imaginary time distances. The fact that renormalized perturbation theory converges is an important fact, based on combinatorial and analytic techniques that take advantage of the fermionic nature of the model (it would not be true for a theory involving bosonic degrees of freedom): it allows us to fully keep track of the effects of the lattice and, more in general, of the irrelevant terms appearing in the original Hamiltonian or generated under the RG flow by the iterative, multiscale integration scheme, which may affect the values of finite quantities, such as the Kubo conductivity.
[…]
Next, we start the integration of the infrared degrees of freedom, which is performed iteratively, by using the decomposition $\psi^{(\leq 0)}_{\omega}$ as $\psi^{\pm\,(\leq 0)}_{\omega,{\bf x}}=\sum_{h<h^{\prime}\leq 0}\psi^{\pm\,(h^{\prime})}_{\omega,{\bf x}}+\psi^{\pm\,(\leq h)}_{\omega,{\bf x}}$, which, for $\omega=+$ (resp. $\omega=$-), is valid for any $h\geq h^{*}_{1}$ (resp. $h\geq h^{*}_{2}$), with $h^{*}_{1}=\lfloor\log_{2}m^{R}_{+}\rfloor$, with $m^{R}_{+}=m^{R}_{-}+6\sqrt{3}t_{2}\sin\phi$, see [33, eq.(4.24)] (resp. $h^{*}_{2}=h^{*}_{2}(m_{-}^{R}):=\lfloor\log_{2}|m^{R}_{-}|\rfloor$, to be interpreted as being -$\infty$ if $m^{R}_{-}=0$). The single-scale Grassmann field $\psi^{(h^{\prime})}_{\omega}$ has propagator that, in momentum space, is supported on momenta $\bf k$$^{\prime}$ such that $|{\bf k}^{\prime}|$ is of the order $2^{h^{\prime}}$, that is, it is bounded from above and below proportionally to $2^{h^{\prime}}$; physically, $\bf k$$^{\prime}$ represents the quasi-momentum, i.e., the crystalline momentum relative to the Fermi point ${\bf k}^{-}_{F}=(0,\vec{k}^{-}_{F})$). We integrate out the fields on scales $0,-1,-2,\ldots$; once we get to scale $h^{*}_{1}$, we fully integrate out the field $\psi^{(\leq h^{*}_{1})}_{+}$, which is massive on that scale, see [33, eq.(4.32) and the lines after]. We then iteratively continue the integration of the fields $\psi^{(h)}_{-}$, with $h^{*}_{2}<h\leq h^{*}_{1}$; if $m^{R}_{-}\neq 0$, once we get to scale $h^{*}_{2}$, we fully integrate out the field $\psi^{(\leq h^{*}_{2})}_{2}$; if $m_{R}^{-}=0$, then $h^{*}_{2}=$-$\infty$ and the iteration continues for infinitely many steps.
After the integration of the higher-scale degrees of freedom, say for definiteness of the scales higher than $h$, with $h^{*}_{2}<h\leq h^{*}_{1}$, the generating function takes the form:
$\mathcal{W}(\varphi,A)$ $=$ $\sum_{h^{\prime}>h}S^{(h^{\prime})}(\varphi,A)$
$+$ $\log\int P(d\psi^{(\leq h)}_{-})e^{-V^{(h)}(\psi^{(\leq h)}_{-})+B^{(h)}(\varphi,A,\psi^{(\leq h)}_{-})},$
where $S^{(h^{\prime})}(\varphi,A)$ is the single-scale contribution to the generating function, while the effective potential $V^{(h)}$ and the effective source term $B^{(h)}$ explicitly depend on the fluctuation field $\psi^{(\leq h)}_{-}$, i.e., $V^{(h)}(0)=B^{(h)}(\varphi,A,0)=0$ (the source term also explicitly depends on the external fields: $B^{(h)}(0,0,\psi^{(\leq h)}_{-})=0$). $P(d\psi^{(\leq h)}_{-})$ is the Grassmann Gaussian integration with propagator that, in Fourier space, reads
$\hat{g}^{(\leq h)}_{-}({\bf k}^{\prime})=\chi_{h}({\bf k}^{\prime})$ (42)
$\ \ \cdot\begin{pmatrix}-iZ_{1,h}k_{0}^{\prime}+m^{R}_{-}&{-\frac{2}{3}v_{h}\Omega^{*}(\vec{k}_{F}^{-}+\vec{k}^{\prime})}\\ {-{\frac{2}{3}v_{h}\Omega(\vec{k}_{F}^{-}+\vec{k}^{\prime})}}&-iZ_{2,h}k_{0}^{\prime}-m^{R}_{-}\end{pmatrix}^{-1}(1+O({\bf k}^{\prime})),$
where: $\chi_{h}({\bf k}^{\prime})\equiv\chi_{0}(2^{-h}{\bf k}^{\prime})$ is a smooth cutoff function supported on quasi-momenta $|{\bf k}^{\prime}|<2^{h}$ ($\chi_{0}$ is a smooth version of the characteristic function of a small ball centered at the origin; we choose it to be a compactly supported $C^{\infty}$-function, supported on the ball of radius $1/3$ centered at the origin, and identically equal to $1$ on the ball of radius $1/4$ centered at the origin); $Z_{1,h},Z_{2,h}$ are the effective wave function renormalizations, and $v_{h}$ is the effective Fermi velocity, which converge exponentially fast to their infrared limits, i.e., there exist $Z_{1}(m^{R}_{-}),Z_{2}(m^{R}_{-}),v(m^{R}_{-})$, analytic in $U$, such that $|Z_{i,h}-Z_{i}(m^{R}_{-})|,|v_{h}-v(m^{R}_{-})|\leq C_{\theta}|U|2^{\theta h}$, for any $0\leq\theta<1$ and some $C_{\theta}>0$, diverging as $\theta\to 1^{-}$; the dependence upon $m^{R}_{-}$ of these effective constants is continuous: $|Z_{i}(m^{R}_{-})-Z_{i}(0)|\leq C_{\theta}|U||m^{R}_{-}|^{\theta}$ and $|v(m^{R}_{-})-v(0)|\leq C_{\theta}|U||m^{R}_{-}|^{\theta}$, for any $0<\theta<1$ and some $C_{\theta}>0$.
Remark 2.
The constants $Z_{1},Z_{2},v$ in the statement of Proposition III.1 are nothing but the values of the corresponding constants at zero mass: $Z_{i}:=Z_{i}(0)$ and $v:=v(0)$. Similarly, $\gamma:=\gamma(0)$, where $\gamma(m^{R}_{-})$ is the mass-dependent effective vertex function defined below, see the discussion after eq.(44).
The function $\Omega$ that appears in the propagator (42) is the same as that in the Bloch Hamiltonian (7). For $|\vec{k}^{\prime}|\ll 1$, one has $\Omega(\vec{k}_{F}^{\omega}+\vec{k}^{\prime})=\frac{3}{2}(ik_{1}^{\prime}+\omega k_{2}^{\prime})+O(|\vec{k}^{\prime}|^{2})$, so in principle in the RHS of (42) we could replace $\Omega$ by its linearization, and include the higher-order terms in the remainder $O(|\bf k^{\prime}|)$. However, we prefer to retain the full lattice function $\Omega(\vec{k}_{F}^{\omega}+\vec{k}^{\prime})$ in order to explicitly preserve the discrete rotational symmetry of [33, eq.(4.7)], which would be violated if $\Omega$ was replaced by its linear approximation.
The effective potential $V^{(h)}$ in (IV.1) can be represented as follows:
$V^{(h)}(\psi)=\!\int\!\tfrac{d{\bf k}^{\prime}}{(2\pi)^{3}}\!\big[2^{h}\xi_{-,h}\hat{\psi}^{+}_{{\bf k}^{\prime}}\hat{\psi}^{-}_{{\bf k}^{\prime}}+2^{h}\delta_{-,h}\hat{\psi}^{+}_{{\bf k}^{\prime}}\sigma_{3}\hat{\psi}^{-}_{{\bf k}^{\prime}}$
$\ +\hat{\psi}^{+}_{{\bf k}^{\prime}}\begin{pmatrix}-iz_{1,-,h}k_{0}^{\prime}&{-{\frac{2}{3}u_{-,h}\Omega^{*}(\vec{k}_{F}^{-}+\vec{k}^{\prime})}}\\ {-{\frac{2}{3}u_{-,h}\Omega(\vec{k}_{F}^{-}+\vec{k}^{\prime})}}&-iz_{2,-,h}k_{0}^{\prime}\end{pmatrix}\hat{\psi}^{-}_{{\bf k}^{\prime}}\big]$
$\ +{\mathcal{R}}V^{(h)}(\psi),$ (43)
where: $\xi_{-,h},\delta_{-,h}$ are two running counterterms, satisfying $|\xi_{-,h}|,\,|\delta_{-,h}|\leq C_{\theta}|U|2^{\theta h}$ for $0\leq\theta<1$ and $C_{\theta}>0$; $z_{1,-,h},z_{2,-,h},u_{-,h}$ are the single-step contributions to the effective couplings $Z_{1,-,h},Z_{2,-,h},v_{h}$, satisfying the same bounds as the running counterterms, i.e., $|z_{1,-,h}|,|z_{2,-,h}|,|u_{-,h}|\leq C_{\theta}|U|2^{\theta h}$; $\mathcal{R}V^{(h)}$ is the irrelevant part of $V^{(h)}$, which is an infinite linear combination of even monomials of the form $\int d{\bf x}_{1}\cdots d{\bf x}_{n}W_{n,p}^{(h)}(\underline{\bf x};\underline{\alpha})D^{\underline{\alpha}}\Psi(\underline{\bf x})$, where ${\bf x}_{i}=(x_{0,i},\vec{x}_{i})$, $\int d{\bf x}_{i}$ is shorthand for $\int_{0}^{\beta}x_{0,i}\sum_{\vec{x}_{i}\in\Lambda}$, $\underline{\alpha}=((p_{1},i_{1}),\ldots,(p_{n},i_{n}))$, $\underline{i}=(i_{1},\ldots,i_{n})$, $\underline{\bf x}=({\bf x},\ldots,{\bf x}_{n})$, and $D^{\underline{\alpha}}\Psi(\underline{\bf x})=\partial^{p_{1}}_{i_{1}}\psi^{+}_{{\bf x}_{1}}\partial^{p_{2}}_{i_{2}}\psi^{-}_{{\bf x}_{2}}\cdots\partial^{p_{n-1}}_{i_{n-1}}\psi^{+}_{{\bf x}_{n-1}}\partial^{p_{n}}_{i_{n}}\psi^{-}_{{\bf x}_{n}}$ with $p_{j}=0,1$ and $i_{j}=0,1,2$: irrelevance of the monomial means that $n$ and $p:=\sum_{j=1}^{n}p_{j}$ are such that the scaling dimension of the monomial
$d_{sc}(n,p)=3-n-p$
is negative; finally, the kernels of the irrelevant terms satisfy $\|W^{(h)}_{n,p}\|_{h}\leq C^{n}|U|^{{\max}\{1,\frac{n}{2}-1\}}2^{d_{sc}(n,p)h}$, where $\|W^{(h)}_{n,p}\|_{h}:=\int d{\bf x}_{2}\cdots d{\bf x}_{n}|W^{(h)}_{n,p}(\underline{\alpha};\underline{\bf x})|e^{\kappa\sqrt{2^{h}\delta(\underline{\bf x})}}$ with $\kappa$ a sufficiently small positive constant and $\delta(\underline{\bf x})$ the tree distance among the elements of $\underline{\bf x}$ (also known as the Steiner diameter of $\{{\bf x}_{1},\ldots,{\bf x}_{n}\}$, see [42, footnote 19]).
Similarly, the effective source term $B^{(h)}$ can be represented as:
$B^{(h)}(\varphi,A,\psi)=(\psi,\varphi)+(J_{h},A)+\mathcal{R}B^{(h)}(\varphi,A,\psi),$ (44)
[…]
$d_{sc}(l,m,n,p):=3-2l-m-n-p;$
finally, the kernels $W_{l,m,n,p}^{(h)}$ satisfy a norm bound analogous to that of the kernels of the effective potential, namely
$\|W_{l,m,n,p}^{(h)}\|_{h}\leq C^{l+m+n}2^{d_{sc}(l,m,n,p)h}.$ (45)
In (IV.1), the single-scale contribution to the generating function, $W^{(h)}$, admits a representation similar to $V^{(h)}$ and $B^{(h)}$, that is, it can be written as an infinite linear combination of monomials of the form $\int d\underline{\bf y}\,d\underline{\bf z}\,W_{l,m,0,0}^{(h)}(\underline{\bf y},\underline{\bf z};\underline{\mu})\varphi(\underline{\bf y})A_{\underline{\mu}}(\underline{\bf z})$, where the kernels $W_{l,m,0,0}^{(h)}$ satisfy the same norm bound as (45), with $n=p=0$.
At each step of the iteration leading to (IV.1), the marginal quadratic terms in $\psi$, i.e., those in the second line of (43), are re-absorbed into the Gaussian Grassmann measure, thus contributing to the iterative dressing of the propagator. Next, the modified propagator is decomposed into a single-scale contribution, associated with momenta on scale $h$, plus an infrared contribution, associated with momenta on scales smaller than $h$; the contribution from the single-scale propagator at scale $h$ is integrated out, and a formula analogous to (IV.1) is obtained, with $h$ replaced by h1h-1.
[…]
$W_{l,m,n,p}^{(h)}=\sum_{\tau\in\mathcal{T}_{l,m,n,p}^{(h)}}W_{l,m,n,p}[\tau],$
where $W_{l,m,n,p}[\tau]$ is the tree value of $\tau$. The tree expansion is a combinatorially convenient (and better behaved) resummation of the more naive expansion in Feynman diagrams: each tree value is nothing but the sum over a family of Feynman diagrams with given vertices (represented by the endpoints of the Gallavotti–Nicolò trees) and given scale labels associated with the propagators, satisfying a hierarchical structure of clusters into clusters compatible with the structure of the tree (a cluster is a Feynman sub-diagram with propagators whose scale labels are all higher than those of the lines exiting from the sub-diagram itself); see [44, Section 5] for a review. In particular, each (rooted) tree $\tau\in\mathcal{T}_{l,m,n,p}^{(h)}$ contributing to $W^{(h)}_{l,m,n,p}$ has root on scale $h$ and endpoints on higher scales, between $h+1$ and $1$; the endpoints on scales lower than $1$ correspond to ‘dressed’ vertices, associated with either one of the two contributions in the first line of (43) or one of the first two contributions on the RHS of (44); the endpoints on scale $1$ correspond to ‘bare’ vertices, associated with one of the contributions of the ‘bare’ action $V(\psi)+(\varphi,\psi)+(j,A)$ in the exponent on the RHS of (37). Note that any endpoint associated with a quartic contribution (the first on the RHS of (39)) is necessarily on scale 1; this convention is related to the irrelevance of the quartic interaction: $d_{sc}(0,0,4,0)=-1$. We shall refer to the endpoints associated with a quartic interaction or to one of the running counterterms (see (43) and following lines) as ‘interaction endpoints’. The following key estimate on the sum of the tree values over trees having at least one interaction endpoint is valid, and will be used in the rest of the proof:
$\sum_{\begin{subarray}{c}\tau\in\mathcal{T}_{l,m,n,p}^{(h)}:\\ N_{\text{int}}(\tau)\geq 1\end{subarray}}\!\!\!\!\!\!\!\|W_{l,m,n,p}[\tau]\|_{h}\leq C_{\theta}^{l+m+n}|U|2^{d_{sc}(l,m,n,p)h}2^{\theta h},$ (46)
where $N_{\text{int}}(\tau)$ is the number of interaction endpoints, $0\leq\theta<1$, and $C_{\theta}$ is a positive constant such that $C_{\theta}\to+\infty$ as $\theta\to 1^{-}$. The good factor $2^{\theta h}$ physically represents a dimensional gain coming from the fact that the quartic interaction is irrelevant in the infrared, with scaling dimension 1-1 (the ‘$1$’ in the condition $\theta<1$ is the opposite of the scaling dimension of the quartic terms).
IV.2 The differentiable and singular contributions
In view of (11), we need to compute the left derivative of the current-current response function $\hat{K}_{12}(p_{0})$ at $p_{0}=0$. Making the dependence upon $m_{-}^{R}$ explicit, from the RG construction we obtain the rewriting:
$\hat{K}_{12}(p_{0};m^{R}_{-})=\sum_{h=h^{*}_{2}}^{0}\hat{W}^{(h)}_{0,2,0,0}((p_{0},\vec{0});(1,2)),$ (47)
where $\hat{W}^{(h)}_{0,2,0,0}(\mathbf{p};(\mu_{1},\mu_{2}))$ is the Fourier transform of the kernel of $W^{(h)}(\varphi,A)$ with $l=0$ and $m=2$ defined after (45). As reviewed above, $\hat{W}^{(h)}_{0,2,0,0}$ can be written as a sum over Gallavotti–Nicolò trees, which induces the following decomposition of the current-current response function:
$\hat{K}_{12}(p_{0};m^{R}_{-})=\hat{K}_{12}^{\text{diff}}(p_{0};m^{R}_{-})+\hat{K}_{12}^{\text{sing}}(p_{0};m^{R}_{-}),$ (48)
where $\hat{K}_{12}^{\text{diff}}(p_{0};m^{R}_{-})$ (resp. $\hat{K}_{12}^{\text{sing}}(p_{0};m^{R}_{-})$) is the sum over the scales $h$ from $h^{*}_{2}$ to 0 of the sum over trees $\tau\in\mathcal{T}_{0,2,0,0}^{(h)}$ with at least one interaction endpoint (resp. with no interaction endpoints) of $\hat{W}_{0,2,0,0}[\tau]((p_{0},0);m^{R}_{-})$. As we shall shortly see, the first contribution on the RHS of (48) is continuously differentiable with respect to $p_{0}$, while the second is ‘singular’, i.e., its derivative with respect to $p_{0}$ is dimensionally logarithmically divergent as $p_{0}\to 0$. As explained below, the decomposition (48) is the basis for the rewriting (31).
Let us first focus on the first term on the RHS of (48): from (46), we find that
$|\hat{K}_{12}^{\text{diff}}(p_{0};m^{R}_{-})|\leq\sum_{h=h^{*}_{2}}^{0}C_{\theta}^{2}|U|2^{h(\theta+1)},$
[…]
$|\partial_{p_{0}}\hat{K}_{12}^{\text{diff}}(p_{0};m^{R}_{-})|\leq\sum_{h=h^{*}_{2}}^{0}C_{\theta}^{2}|U|2^{h\theta},$
[…]
$|\partial_{p_{0}}\hat{K}_{12}^{\text{diff}}(p_{0};m^{R}_{-})-\partial_{p_{0}}\hat{K}_{12}^{\text{diff}}(p_{0};0)|$ (49)
$\qquad\leq|m_{-}^{R}|\sum_{h=h^{*}_{2}}^{0}C_{\theta}^{2}|U|2^{h(\theta-1)}+\sum_{h<h^{*}_{2}}C_{\theta}^{2}|U|2^{h\theta}$
$\qquad=C^{\prime}_{\theta}|U||m_{-}^{R}|2^{h^{*}_{2}(\theta-1)}+C^{\prime\prime}_{\theta}|U|2^{h^{*}_{2}\theta}=C^{\prime\prime\prime}_{\theta}|U||m_{-}^{R}|^{\theta},$
where in the last identity we used the definition of $h^{*}_{2}:=\lfloor\log_{2}|m_{-}^{R}|\rfloor$.
Let us now consider the ‘singular’ term $\hat{K}_{12}^{\text{sing}}(p_{0};m^{R}_{-})$, which is by definition the contribution to the RHS of (47) from trees with no interaction endpoints and two endpoints associated with source terms $(J_{h},A)$ and $(J_{h^{\prime}},A)$, with $h,h^{\prime}\geq h^{*}_{2}$. This contribution is very explicit: it is just the sum over scales of the dressed ‘bubble diagram’:
$\mathbf{p}$$\mathbf{k}^{\prime}+\mathbf{p}$$\mathbf{k}^{\prime}$$\mathbf{p}$$1,h^{\prime}$$2,h$
with $\mathbf{p}=(p_{0},\vec{0})$, whose explicit expression is
$\begin{split}&-\int\frac{d\mathbf{k}^{\prime}}{2\pi|\mathcal{B}|}\mbox{Tr}\Big\{\Gamma_{h\vee h^{\prime},1}(\vec{k}^{\prime},\vec{p})\hat{g}^{(h)}_{m_{-}^{R}}(\mathbf{k}^{\prime})\\ &\qquad\qquad\qquad\cdot\Gamma_{h\vee h^{\prime},2}(\vec{k}^{\prime}+\vec{p},-\mathbf{p})\hat{g}^{(h^{\prime})}_{m_{-}^{R}}(\mathbf{k}^{\prime}+\mathbf{p})\Big\},\end{split}$ (50)
where $h\vee h^{\prime}:=\max\{h,h^{\prime}\}$, the functions $\Gamma_{h,\mu}(\vec{k}^{\prime},\vec{p})$ are the dressed vertex functions whose properties are spelled out after (44), and, if $h>h^{*}_{2}(m^{R}_{-})$, $\hat{g}^{(h)}_{m_{-}^{R}}(\mathbf{k}^{\prime})$ is the single-scale counterpart of (42), satisfying
$\hat{g}^{(h)}_{m_{-}^{R}}({\bf k}^{\prime})=f_{h}({\bf k}^{\prime})$ (51)
$\ \ \cdot\begin{pmatrix}-iZ_{1,h}k_{0}^{\prime}+m^{R}_{-}&v_{h}(ik_{1}^{\prime}+k_{2}^{\prime})\\ v_{h}(-ik_{1}^{\prime}+k_{2}^{\prime})&-iZ_{2,h}k_{0}^{\prime}-m^{R}_{-}\end{pmatrix}^{-1}(1+O({\bf k}^{\prime})),$
where $f_{h}(\mathbf{k}^{\prime}):=\chi_{h}(\mathbf{k}^{\prime})-\chi_{h-1}(\mathbf{k}^{\prime})$ is supported on momenta of order $2^{h}$, i.e., in the region $\tfrac{2^{h}}{8}\leq|\mathbf{k}^{\prime}|\leq\tfrac{2^{h}}{3}$ (if $h=h^{*}_{2}(m^{R}_{-})$ we shall identify $\hat{g}^{(h)}_{m_{-}^{R}}(\mathbf{k}^{\prime})$ with $\hat{g}^{(\leq h)}_{-}(\mathbf{k}^{\prime})$ in (42)). Moreover, the effective constants $Z_{1,h},Z_{2,h},v_{h}$ satisfy the properties described in the paragraph after (42). For later reference, we note that, on their supports (which are the same as the support of $f_{h}$ or of $\chi_{h}$, depending on whether $h>h^{*}_{2}(m^{R}_{-})$ or $h=h^{*}_{2}(m^{R}_{-})$), the derivatives of $g^{(h)}_{m_{-}^{R}}(\mathbf{k}^{\prime})$ with respect $k_{0}$ and/or $m_{-}^{R}$ satisfy the following dimensional bounds:
$\begin{split}&\big|\partial_{k_{0}^{\prime}}\hat{g}^{(h)}_{m_{-}^{R}}({\bf k^{\prime}})\big|,\ \big|\partial_{m_{-}^{R}}\hat{g}^{(h)}_{m_{-}^{R}}({\bf k^{\prime}})\big|\leq C2^{-2h},\\ &\big|\partial_{k_{0}^{\prime}}\partial_{m_{-}^{R}}\hat{g}^{(h)}_{m_{-}^{R}}({\bf k}^{\prime})\big|\leq C2^{-3h}.\end{split}$ (52)
Let us go back to the decomposition (48): plugging it into (11), we find that the transverse conductivity of the reference model can be rewritten as
$\sigma_{12}^{R}(U,m_{-}^{R},\phi)=\sigma_{12}^{\text{diff}}(U,m_{-}^{R},\phi)+\sigma_{12}^{\text{sing}}(U,m_{-}^{R},\phi),$ (53)
where we defined
$\begin{split}\sigma_{12}^{\text{diff}}(U,m_{-}^{R},\phi)&=\frac{1}{|\vec{\ell}_{1}\times\vec{\ell}_{2}|}\partial_{p_{0}}\hat{K}_{12}^{\text{diff}}(0;m_{-}^{R}),\\ \sigma_{12}^{\text{sing}}(U,m_{-}^{R},\phi)&=\frac{1}{|\vec{\ell}_{1}\times\vec{\ell}_{2}|}\partial_{p_{0}}^{-}\hat{K}_{12}^{\text{sing}}(0;m_{-}^{R}).\end{split}$ (54)
In particular, from the explicit expression of $\hat{K}_{12}^{\text{sing}}$, we find that
$\sigma_{12}^{\text{sing}}(U,m_{-}^{R},\phi)=-\!\!\sum_{h,h^{\prime}=h^{*}_{2}}^{0}\!\!\partial^{-}_{p_{0}}\int\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m^{R}_{-}),$ (55)
with $h^{*}_{2}=h^{*}_{2}(m_{-}^{R})$ and
$\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m^{R}_{-})$ $=$ $\mbox{Tr}\Big\{\Gamma_{h\vee h^{\prime},1}(\vec{k}^{\prime},\vec{p})\hat{g}^{(h)}_{m_{-}^{R}}(\mathbf{k}^{\prime})$
$\cdot$ $\Gamma_{h\vee h^{\prime},2}(\vec{k}^{\prime}+\vec{p},-\vec{p})\hat{g}^{(h^{\prime})}_{m_{-}^{R}}(\mathbf{k}^{\prime}+\mathbf{p})\Big\}.$
IV.3 Completion of the proof
We now use the decomposition (53) and the bounds derived above to prove the continuity property (34). By definition,
$\sigma_{12}^{R,(2)}(U,m^{R}_{-},\phi)$ $=$ $\sigma_{12}^{\text{diff}}(U,m_{-}^{R},\phi)$
$+$ $\big[\sigma_{12}^{\text{sing}}(U,m_{-}^{R},\phi)-\sigma_{12}^{R,(1)}(U,m^{R}_{-},\phi)\big],$
with $\sigma^{R,(1)}_{12}$ as in (32). Using the definition of $\sigma_{12}^{\text{diff}}$ (see the first line of (54)), and the Hölder continuity property (49), we immediately find that $\lim_{m_{-}^{R}\to 0}\sigma_{12}^{\text{diff}}(U,m_{-}^{R},\phi)=\sigma_{12}^{\text{diff}}(U,0,\phi)$. We are then left with proving that, for $m=\pm\epsilon$,
$\begin{split}&\lim_{\epsilon\to 0^{+}}\big[\sigma_{12}^{\text{sing}}(U,m,\phi)-\sigma_{12}^{R,(1)}(U,m,\phi)\\ &\qquad\ \ -\sigma_{12}^{\text{sing}}(U,0,\phi)+\sigma_{12}^{R,(1)}(U,0,\phi)\big]=0.\end{split}$ (58)
For $m=\pm\epsilon$, recalling that $\mathcal{A}_{\epsilon}$ is the ball of radius $\sqrt{\epsilon}$ centered at the origin, we rewrite $\sigma_{12}^{R,(1)}(U,m,\phi)$ in (32) as
$\sigma_{12}^{R,(1)}(U,m,\phi)\!=-\partial^{-}_{p_{0}}\!\!\!\sum_{h,h^{\prime}=h_{1,\epsilon}}^{h_{2,\epsilon}}\int\limits_{\mathcal{A}_{\epsilon}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)\Big|_{\mathbf{p}=\mathbf{0}},$ (59)
where $h_{1,\epsilon}\equiv h^{*}_{2}(\epsilon)=\lfloor\log_{2}\epsilon\rfloor$, $h_{2,\epsilon}=\lfloor\log_{2}(9\sqrt{\epsilon})\rfloor$, $\mathbf{p}=(p_{0},\vec{0})$, and, if $h,h^{\prime}>h^{*}_{2}(m^{R}_{-})$,
$\begin{split}\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)&=-\gamma^{2}f_{h}(\mathbf{k}^{\prime})f_{h^{\prime}}(\mathbf{k}^{\prime}+\mathbf{p})\\ &\cdot\textup{Tr}\left\{\sigma_{2}S_{m}(\mathbf{k}^{\prime})\sigma_{1}{S}_{m}(\mathbf{k}^{\prime}+\mathbf{p})\right\},\end{split}$ (60)
with $\gamma$ and $S_{m}({\bf k})$ the same as in (32), and where $f_{h}$ was defined after (51) (if $h=h^{*}_{2}(m^{R}_{-})$, then $f_{h}$ in (60) should be replaced by $\chi_{h}$, and similarly for $h^{\prime}$). Note that the upper bound in the sum over $h,h^{\prime}$ in (59) is due to the fact that, for $\mathbf{p}$ small enough, the support of $\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)$ intersects $\mathcal{A}_{\epsilon}$ only if $h,h^{\prime}\leq h_{2,\epsilon}$. Note also that $\sigma_{12}^{R,(1)}(U,0,\phi)$ admits a decomposition similar to (59), with $h_{1,\epsilon}$ replaced by -$\infty$ and $\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)$ replaced by $\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)$.
We now plug the decompositions (55) and (59), as well as the analogous one for $\sigma_{12}^{R,(1)}(U,0,\phi)$, in the expression in brackets in (58), and rewrite it as the sum of three terms, $I_{1}$, $I_{2}$, $I_{3}$, which will separately be shown to tend to zero as $\epsilon\to 0^{+}$. These are defined as follows (recall that $\mathbf{p}=(p_{0},\vec{0})$, that $m=\pm\epsilon$, and let $h_{3,\epsilon}=\lfloor\log_{2}\sqrt{\epsilon}\rfloor$):
$\begin{split}&\thinspace I_{1}\!=\!\!\!\!\sum_{h,h^{\prime}=h_{3,\epsilon}}^{0}\int\limits_{\mathcal{A}_{\epsilon}^{c}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\partial_{p_{0}}\Big[\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)-\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)\Big]\Big|_{\mathbf{p}=\mathbf{0}}\\ &\thinspace I_{2}\!=\!\!\!\!\sum_{h,h^{\prime}=h_{1,\epsilon}}^{h_{2,\epsilon}}\int\limits_{\mathcal{A}_{\epsilon}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\partial_{p_{0}}\Big[\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)-\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)\Big]\Big|_{\mathbf{p}=\mathbf{0}}\\ &\thinspace I_{3}\!=\partial^{-}_{p_{0}}\!\!\sum_{h,h^{\prime}=-\infty}^{h_{2,\epsilon}}\int\limits_{\mathcal{A}_{\epsilon}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\Big[\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)-\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)\Big]\Big|_{\mathbf{p}=\mathbf{0}}\end{split}$ (61)
In the definition of $I_{1}$, the lower bound on the sum over $h,h^{\prime}$ is due to the fact that, for $\mathbf{p}$ small enough, the support of $\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};m)$ intersects the complement of $\mathcal{A}_{\epsilon}$ only if $h,h^{\prime}\geq h_{3,\epsilon}$). Note that in the definitions of $I_{1}$ and $I_{2}$, we could move the left derivatives inside the sums, because they run over finitely many terms; moreover, each term in each of the two sums is smooth in $\mathbf{p}$ and absolutely integrable and, therefore, in both cases we could change the left derivatives to full derivatives and move them inside the integral. Note also that, without loss of generality, the sums over $h,h^{\prime}$ in the definitions of $I_{1},I_{2}$ can be restricted to pairs of scale labels such that $|h-h^{\prime}|\leq 1$: in fact, if $|h-h^{\prime}|\geq 2$ and $\mathbf{p}$ is small enough, the summands are zero, due to the compact support properties of $\mathcal{F}_{h,h^{\prime}}$ and $\mathcal{G}_{h,h^{\prime}}$.
[…]
$|I_{1}|\leq C\epsilon\sum_{h\geq h_{3,\epsilon}}2^{-h}=C^{\prime}\sqrt{\epsilon},$ (62)
where we used the definition of $h_{3,\epsilon}=\lfloor\log_{2}\sqrt{\epsilon}\rfloor$ in the identity.
[…]
$|I_{2}|\leq C\sum_{h=h_{1,\epsilon}}^{h_{2,\epsilon}}(2^{h}+|U|2^{\theta h})\leq C^{\prime}(\sqrt{\epsilon}+|U|\epsilon^{\theta/2}),$ (63)
where we used the definition of $h_{2,\epsilon}={\lfloor\log_{2}(9\sqrt{\epsilon})\rfloor}$ in the second bound.
Finally, consider $I_{3}$. This term is a bit more delicate, because the number of terms involved in the sum is infinite and, therefore, we cannot a priori exchange the left derivative with the sum and then with the integral. We thus fix a $p_{0}<0$ small enough and derive a bound on the incremental ratio
$\Delta_{p_{0}}\sum_{h,h^{\prime}=-\infty}^{h_{2,\epsilon}}\int\limits_{\mathcal{A}_{\epsilon}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\Big[\mathcal{G}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)-\mathcal{F}_{h,h^{\prime}}(\mathbf{k}^{\prime},\mathbf{p};0)\Big],$ (64)
where the action of $\Delta_{p_{0}}$ on a function $F(p_{0})$ (recall that $\mathbf{p}=(p_{0},\vec{0})$) is defined as $\Delta_{p_{0}}F(p_{0}):=\frac{1}{p_{0}}\big[F(p_{0})-F(0)\big]$. In (64) we first focus on the contribution from the diagonal terms in the sum, i.e., those with $h^{\prime}=h$. These can be written as:
$\begin{split}&\sum_{h=-\infty}^{h_{2,\epsilon}}\int\limits_{\mathcal{A}_{\epsilon}}\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\int_{0}^{1}\!\!ds\,\mbox{Tr}\Big\{\Gamma_{h,1}(\vec{k}^{\prime},\vec{0})\hat{g}^{(h)}_{0}(\mathbf{k}^{\prime})\Gamma_{h,2}(\vec{k}^{\prime},\vec{0})\\ &\ \cdot\partial_{p_{0}}\hat{g}^{(h)}_{0}(\mathbf{k}^{\prime}+s\mathbf{p})+\gamma^{2}\sigma_{2}S_{0}^{(h)}(\mathbf{k}^{\prime})\sigma_{1}\partial_{p_{0}}S_{0}^{(h)}(\mathbf{k}^{\prime}+s\mathbf{p})\Big\},\end{split}$ (65)
[…]
where we used the definition of $h_{2,\epsilon}$ in the second bound.
We next focus on the contributions from the off-diagonal terms in (64), i.e., those with $h^{\prime}\neq h$, such that either $\max\{h,h^{\prime}\}<h_{2,\epsilon}$, or $\max\{h,h^{\prime}\}=h_{2,\epsilon}$ and $\min\{h,h^{\prime}\}\leq h_{2,\epsilon}-2$ (note that for these terms, if $p_{0}$ is small enough, ${\text{supp}}(f_{h})\cap{\text{supp}}(f_{h^{\prime}}(\cdot+{\bf p}))$ is strictly contained in $\mathcal{A}_{\epsilon}$). After symmetrization under the exchange of $h$ with $h^{\prime}$, these can be rewritten as
$\thinspace\Big(\sum_{h^{\prime}<h<h_{2,\epsilon}}\!\!\!+\sum_{\begin{subarray}{c}h^{\prime}\leq h-2:\\ h=h_{2,\epsilon}\end{subarray}}\Big)\int\!\frac{d\mathbf{k}^{\prime}}{(2\pi)^{3}}\Big(I_{h^{\prime},h}^{1,2}(\mathbf{k}^{\prime},p_{0})\ -I^{2,1}_{h^{\prime},h}(\mathbf{k}^{\prime},-p_{0})\Big),$ (67)
where, recalling that $\mathbf{p}=(p_{0},\vec{0})$,
$\begin{split}&I_{h^{\prime},h}^{i,j}(\mathbf{k}^{\prime},p_{0})=\int_{0}^{1}\!\!ds\,\mbox{Tr}\Big\{\Gamma_{h,i}(\vec{k}^{\prime},\vec{0})\hat{g}^{(h^{\prime})}_{0}(\mathbf{k}^{\prime})\Gamma_{h,j}(\vec{k}^{\prime},\vec{0})\\ &\ \cdot\partial_{p_{0}}\hat{g}^{(h)}_{0}(\mathbf{k}^{\prime}+s\mathbf{p})+\gamma^{2}\sigma_{j}S_{0}^{(h^{\prime})}(\mathbf{k}^{\prime})\sigma_{i}\partial_{p_{0}}S_{0}^{(h)}(\mathbf{k}^{\prime}+s\mathbf{p})\Big\}.\end{split}$
By the same considerations spelled out after (65), we find that, for $h^{\prime}<h$ and $(i,j)=(1,2),(2,1)$,
$\int d\mathbf{k^{\prime}}\big|I^{i,j}_{h^{\prime},h}(\mathbf{k}^{\prime},p_{0})\big|\leq({\rm const.})2^{3h^{\prime}}\cdot 2^{-h^{\prime}}\cdot 2^{-2h}\cdot(2^{h}+|U|2^{\theta h}),$
so that
[…]
In conclusion, $|I_{3}|\leq C^{\prime}(\sqrt{\epsilon}+|U|\epsilon^{\theta/2})$ and, therefore, the three terms $I_{1},I_{2},I_{3}$ all tend to zero as $\epsilon\to 0^{+}$. This proves (58), thus concluding the proof of the continuity bound (34), as desired.
V Conclusion
We studied the topological phase transition in an interacting version of the Haldane model, and proved that on the critical lines separating the normal and the topological insulating phases the transverse conductivity is quantized in semi-integer values, equal to the average of the integer values on either side of critical line. Together with the quantization results regarding the critical longitudinal conductivity proved in [32] and the off-critical transversal conductivity proved in [39, 33], our result proves the universality of the conductivity matrix in the whole topological phase diagram of the interacting Haldane model, provided the interaction strength is small enough compared to the bandwidth. Even though, for definiteness, we focus on a specific class of interacting lattice systems in this paper, we believe that our result applies more generally to the transition between distinct Hall phases of lattice interacting electron systems characterized at criticality by a semimetallic behavior, in the absence of on-site disorder.
There are several open problems related to the results and methods introduced in this work, which deserve to be investigated. First, it would be interesting to consider the critical interacting Haldane model in a domain with boundary (say, in the half-plane) and investigate the nature of the edge theory: would it be possible to define and compute an edge Hall coefficient in such a critical, semimetallic, regime, matching the bulk value computed in this paper?
Similar questions, concerning both the bulk and edge transport coefficients, and the construction of a topological phase diagram analogous to the one in Fig.2, can be asked for the thermal Hall conductance in the small temperature limit, see [45], where a Kubo-like formula for the bulk thermal conductance is proposed and related to its edge counterpart, which is supposedly given by the boundary chiral central charge.
Finally, it would be very nice to prove the universality of the longitudinal and transverse conductivities of critical (or quasi-critical) Hall systems in the presence of weak, marginally irrelevant, on-site disorder, in the spirit of [36], see also [46, 47, 48]. Marginally irrelevant contributions may in principle affect finite quantities at criticality such as the longitudinal and transverse conductivities (this is already the case for irrelevant contributions, such as those induced by lattice and short-range interaction effects, as discussed in this paper, and even more for marginally irrelevant ones): therefore, their effect deserves to be studied beyond the existing approximate RG or mean-field schemes. While a non-perturbative proof of universality in the disordered case seems beyond reach of the available rigorous methods, a systematic proof at all orders in renormalized perturbation theory may be within reach. We hope to come back to this and the aforementioned problems in future publications.
Acknowledgements.
We thank Eduardo Fradkin, Marcello Porta, and Roberto Raimondi for useful comments, references and discussions. We gratefully acknowledge financial support from: the European Research Council through the ERC CoG UniCoSM, grant agreement n. 724939 (A.G. and R.R.) and the ERC-StG MaMBoQ, grant agreement n. 802901 (S.F.); MUR, PRIN 2022 project MaIQuFi cod. 20223J85K3 (A.G.); GNFM-INdAM Gruppo Nazionale per la Fisica Matematica.